### v1

# Stationary measures and random walks on $\tilde{A}_{2}$-buildings

Abstract.
Consider a non-elementary group action $G\curvearrowright X$ of a locally compact second countable group $G$ on a separable complete building $X$ of type $\tilde{A}_{2}$. We study $G$-equivariant measurable maps $B\to\text{ch}(X^{\infty})$ between a given $G$-boundary $(B,\nu_{B})$ and the set of chambers of the spherical building at infinity. Using techniques from boundary theory, we prove that such a boundary map exists and is unique. As a consequence, we prove that if $\mu$ is an admissible probability measure on $G$, there is a unique $\mu$-stationary measure $\nu$ supported on the chambers at infinity of the spherical building at infinity $X^{\infty}$. We use this result to study random walks induced by the $G$-action, and we prove that for any base-point $o\in X$, $(Z_{n}o)$ converges almost surely to a regular point of the boundary under a second moment assumption. In the algebraic setting, it implies the following. Let $\Gamma\to\mathbf{G}$ be a Zariski-dense unbounded representation in a reductive $K$-algebraic group with Weyl group of type $A_{2}$, for $K$ a complete field, and let $B(\Gamma,\mu)$ be the Poisson-Furstenberg boundary of $\Gamma$ with respect to an admissible measure $\mu$. Then there is a unique measurable $G$-map $B\to\mathbf{G}(K)/\mathbf{P}(K)$, for $\mathbf{P}$ a minimal parabolic subgroup of $\mathbf{G}$, and there is a unique $\mu$-stationary measure on $\mathbf{G}(K)/\mathbf{P}(K)$, extending results of Furstenberg. Moreover, if $\mu$ is symmetric with finite second moment, the Lyapunov spectrum of the random walk is simple.
1. Introduction
1.1. Random walks and buildings
Buildings were introduced by J. Tits in the early 50s in order to realize exceptional Lie groups as symmetry groups of natural geometric objects. They have proven to be extremely useful in a number of areas linked to the theory of semisimple (or more generally, reductive) algebraic groups. Associated to any (quasi-split) reductive algebraic $K$-group $\mathbf{G}$, there exists a canonical spherical building $\Delta(\mathbf{G})$ on which $G$ acts strongly transitively. If moreover the field is non-Archimedean, then this structure can be refined and gives rise to an affine building called the Bruhat-Tits building $\Delta^{\text{BT}}(\mathbf{G})$ on which $G$ acts, and for which $\Delta(\mathbf{G})$ can be seen as a boundary. In both case, their rich combinatorial and geometric structure gives an effective way to parametrize subgroups of $\mathbf{G}(K)$ such as parabolic (in the spherical case) or parahoric (in the affine case) subgroups. For instance, the set of simplices of $\Delta(\mathbf{G})$ (of a given type) is of the form $\mathbf{G}(K)/\mathbf{Q}(K)$, for $\mathbf{Q}$ a parabolic subgroup, and one can relate bounded subgroups of $\mathbf{G}(K)$ to stabilizers of bounded subsets in $\Delta^{\text{BT}}(\mathbf{G})$.
From a geometric point of view, symmetric spaces of non-compact type and Euclidean buildings are among the most prominent examples of CAT(0) spaces. Tits showed [Tit86] that in dimension larger that 3, any irreducible affine building is the Bruhat-Tits building of an algebraic group over a field with valuation. Dimension-1 affine buildings correspond to trees or, more generally, $\mathbb{R}$-trees. In dimension 2 there exist irreducible affine buildings that are not Bruhat-Tits: they are said exotic. The simplest and most studied family of exotic buildings are those of type $\tilde{A}_{2}$: Ronan showed in a constructive way that there is a uncountable family of exotic buildings of type $\tilde{A}_{2}$ [Ron86]. Discrete buildings of type $\tilde{A}_{2}$ correspond to buildings in which the apartments are tiled by equilateral triangles, and the Bruhat-Tits buildings associated to $\text{SL}(3,K)$, for $K$ a discretely valued field such as $\mathbb{Q}_{p}$, is an example of an affine building of type $\tilde{A}_{2}$.
Notice that many structural results are known for the higher rank simple Lie group $\text{SL}(3,\mathbb{Q}_{p})$: Margulis super-rigidity and the Normal Subgroup Theorem are emblematic examples. It is natural to try and extend these results to groups that act by automorphisms on buildings of type $\tilde{A}_{2}$. This field of research has known much progress recently. For instance it was proven in [LdlSW23] that $\tilde{A}_{2}$-lattices have strong property (T), and in [BFL23] that such groups also satisfy a normal subgroup theorem. This program implies the translation of methods and results on algebraic groups to a purely geometric language, and such a translation is interesting and fruitful in itself. In this paper, we consider a non-elementary action $G\curvearrowright X$ on a $\tilde{A}_{2}$-affine building, and we investigate questions related to measurable boundary maps and random walks on $X$, in the spirit of Furstenberg’s work for linear groups [Fur63, Fur73].
Let $G$ be a discrete group, and let $\mu$ be a probability measure on $G$. Let $(\Omega,\mathbb{P})$ be the probability space $(G^{\mathbb{N}},\mu^{\otimes\mathbb{N}})$, with the product $\sigma$-algebra. The space $\Omega$ is called the space of (forward) increments. We denote by $S:\Omega\rightarrow\Omega$ the shift $S((\omega_{i})_{i})=(\omega_{i+1})_{i}$, and by $f:\omega\in\Omega\mapsto\omega_{1}\in G$ the projection to the first coordinate. The application
$(\omega,n)\in\Omega\times\mathbb{N}\mapsto Z_{n}(\omega)=f(\omega)f(S\omega)\dots f(S^{n-1}\omega)=\omega_{1}\omega_{2}\dots\omega_{n},$
defines the random walk on $G$ generated by the measure $\mu$. Now let $(X,d)$ be a metric space endowed with its Borel $\sigma$-algebra and consider an isometric $G$-action $\phi:G\to\operatorname{Isom}(X)$. The data $(\Omega,\mathbb{P},S,f,\phi)$ defines a random dynamical system
$\Omega\times\mathbb{N}\times X\to X$
$(\omega,n,o)\mapsto\phi(Z_{n}(\omega))o.$
[…]
A key feature of Furstenberg’s approach to these dynamical systems is the use of boundary theory, and the study of stationary measures, i.e. measures that are “invariant” under the dynamical system. A probability measure $\nu$ on $X$ is said to be $\mu$-stationary if it satisfies
$\nu=\mu\ast\nu,$
where $\mu\ast\nu$ is the convolution product between $\mu$ and $\nu$: for all $f$ bounded measurable function on $X$,
[…]
Finding stationary measures on $X$ or on a natural compactification of $X$ and studying their properties is one of the major tasks in the theory of random dynamical systems.
The goal of this paper is to undertake this when $X$ is an affine building of type $\tilde{A}_{2}$, endowed with a non-elementary isometric action by a measured group $(G,\mu)$. Having in mind the deep connections between buildings and algebraic groups, studying stationary measures and random processes on $\tilde{A}_{2}$-buildings can be seen as an extension of the theory of random products in linear groups.
In this paper, we do not restrict ourselves to simplicial buildings, as we work in the context of non-discrete $\tilde{A}_{2}$-buildings. Non-discrete affine buildings are interesting for various reasons. As B. Kleiner and B. Leeb showed in [KL97, Theorem 5.2.1], non-discrete affine buildings arise naturally as asymptotic cones of symmetric spaces of non-compact type and of (even locally finite) affine buildings. The use of non-discrete affine buildings is crucial to their proof of quasi-isometric rigidity of symmetric spaces of non-compact type and affine buildings. Another source of interesting examples is that of (reductive) algebraic groups $\mathbb{G}(K)$ over fields with a dense valuation $v:K\to\mathbb{R}$. Actually, F. Bruhat and J. Tits use root groups $U_{\alpha}$ in order to associate affine buildings to any (quasi-split) reductive group over a field $K$ with valuation in this level of generality [BT84]. As A. Parreau showed in [Par00, Section 3], we can also naturally associate a non-discrete building of type $\tilde{A}_{n-1}$ to $\text{GL}_{n}(K)$ using homothety classes of ultrametric norms. We can find other constructions over Henselian fields with valuation in [KP23] and with more general valuations in ordered abelian groups in [HIL24]. Last, actions on non-discrete buildings appear on the boundary of the real spectrum compactification of character varieties [BIPP21]. Such sources of examples, as well as the possibility to extend previously-known results about discrete buildings, indicate that it is worth studying affine buildings in this more general context.
Affine buildings are CAT(0) spaces. As such, there is a natural bordification of $X$ which is given by the visual bordification $X\cup\partial_{\infty}X$. The space $\partial_{\infty}X$ is the set of equivalence classes of rays, where two rays are equivalent if they are at finite Hausdorff distance. Notice that since $X$ is not assumed locally compact, the space $\partial_{\infty}X$ need not be compact. We can refine this structure to associate a spherical building at infinity $X^{\infty}$, which is a simplicial complex whose simplices of maximal dimension form the (uncountable) set of chambers at infinity, and which we denote by $\text{ch}(X^{\infty})$. For $\mathbf{G}$ a simple algebraic group defined over a non-Archimedean field $K$, the set of chambers at infinity of $\Delta^{BT}(\mathbf{G})$ is identified with the projective variety $\mathbf{G}(K)/\mathbf{P}(K)$, for $\mathbf{P}$ a minimal parabolic subgroup.
1.2. Statements of the results
In the seminal article [Fur63], H. Furstenberg defines the Poisson-Furstenberg boundary $B(G,\mu)$ associated to an (admissible) probability measure $\mu$ on a semisimple group $G$: it is a Borel probability space that represents the future asymptotic directions of the $\mu$-random walk on $G$. Among other features, $(B(G,\mu),\nu_{B})$ represents all the bounded $\mu$-harmonic functions on $G$ in a minimal way. The Poisson-Furstenberg boundary is naturally endowed with a non-singular $G$-action which is amenable in Zimmer’s sense (see [Zim78]) and satisfies a number of interesting ergodic properties, see for instance [Kai03] and [BF14].
A fundamental result in boundary theory is that to every $\mu$-stationary measure $\nu$ on a $G$-space $X$, one can associate a $G$-equivariant measurable map
$\psi:b\in B\rightarrow\nu_{b}\in\text{Prob}(X).$
Conversely, such a map gives rise to a stationary measure defined by
[…]
In particular, if we manage to construct and understand $G$-equivariant measurable maps
$\psi:B(G,\mu)\to\text{ch}(X^{\infty}),$
we obtain a $\mu$-stationary measure $\nu$ on $X$ defined by
[…]
In order to study such $G$-maps, we use tools from boundary theory, and in particular the ergodic properties of the Poisson-Furstenberg boundary. The notion of $G$-boundaries was introduced by U. Bader and A. Furman [BF14] after ideas of Kaimanovich [Kai03]: they are Lebesgue spaces endowed with a Zimmer-amenable action of $G$ satisfying strong ergodic properties. An emblematic example of $G$-boundary is actually the Poisson-Furstenberg boundary associated to an admissible probability measure on $G$, but we study measurable $G$-maps $B\to\text{ch}(X^{\infty})$ for any $G$-boundary.
For the rest of the text, we assume that the action of $G$ on the building $X$ is non-elementary in the following sense: there is no bounded orbit in $X$, nor finite orbit on the boundary (that is, no $G$-invariant geodesic line nor apartment). This assumption is central to our argument. We also make the implicit assumption that the action on the building is by automorphisms in the sense of, for instance, [Rou23, §2.1.13], that is, isometries that preserve the structure of walls. The group of automorphisms contains a finite-index subgroup formed by automorphisms that are type-preserving, see [Rou23, §2.4.6.1]. Hereafter, we therefore always implicitly make the mild assumption that the $G$-action is type-preserving. Moreover, by [SST20, Lemma 3.5], we may assume that $X$ is an $\mathbb{R}$-building in which each vertex is special.
Theorem 1.1.
Let $X$ be a separable complete building of type $\tilde{A}_{2}$ and let $G$ be a locally compact second countable group. Let $G\curvearrowright X$ be a type-preserving and non-elementary action by isometries. Let $(B,\nu)$ be a $G$-boundary. Then there exists a unique measurable map
$B\rightarrow\text{ch}(X^{\infty})$
which is $G$-equivariant.
The separability condition is so that the boundary $\partial_{\infty}X$ is metrizable, which is important for measure-theoretic purposes. Nevertheless, we can get rid of this assumption using a reduction process explained in Section 1.3 below.
Thanks to the duality between $G$-maps $B(G,\mu)\to\text{ch}(X^{\infty})$ and $\mu$-stationary measures, we obtain our main result.
Theorem 1.2.
Let $X$ be a separable complete building of type $\tilde{A}_{2}$ and let $G$ be a locally compact second countable group. Let $G\curvearrowright X$ be a type-preserving and non-elementary action by isometries, and let $\mu$ be an admissible symmetric measure on $G$. Then there exists a unique $\mu$-stationary measure $\nu$ on $\text{ch}(X^{\infty})$. Moreover, we have the decomposition
[…]
where $\psi$ is the unique boundary map given by Theorem 1.1.
Furstenberg proved in [Fur63] that if $G$ is a connected non-compact semisimple real Lie group with finite center, $P\leq G$ is a minimal parabolic subgroup, and $\mu$ is an admissible measure on $G$, then there exists a unique $\mu$-stationary measure on $G/P$. Theorem 1.2 is the analogue of this fact in the non-discrete, possibly exotic case.
Two chambers at infinity are said to be opposite if they belong to the boundary of a unique apartment. We prove that for the stationary measure $\nu$ on $\text{ch}(X^{\infty})$, almost every pair of chambers at infinity are opposite.
Proposition 1.3.
Let $\nu$ be the unique $\mu$-stationary measure on $\text{ch}(X^{\infty})$ given by Theorem 1.2. Then $\nu\otimes\nu$-almost every pair of chambers in $\text{ch}(X^{\infty})$ are opposite.
In other words, the measurable equivariant map $\psi:B\to\text{ch}(X^{\infty})$ is almost surely antipodal. Such a statement can be useful in a number of applications. In particular, it is needed for the study of the limit laws of the random walk $(Z_{n}o)$ generated by $\mu$. The following result states that this random walk converges almost surely to a chamber at infinity. We say that the measure $\mu$ has finite second moment if
[…]
Let $\mathfrak{a}^{+}$ be the fundamental Weyl chamber, and denote by $\mathfrak{a}^{++}$ its interior. For $\lambda\in\mathfrak{a}^{+}$, we say that that a sequence $(x_{n})$ in $X$ is $\lambda$-regular if it sublinearly tracks a $\lambda$-ray, see Section 7.1 for precise definitions.
Theorem 1.4 (Simplicity of the Lyapunov spectrum).
Let $X$ be a separable complete building of type $\tilde{A}_{2}$, $G$ a locally compact second countable group and $G\curvearrowright X$ a non-elementary action on $X$ by isometries. Let $\mu$ be a symmetric and admissible probability measure with finite second moment. Then there exists a regular vector $\lambda\in\mathfrak{a}^{++}$ such that for almost every $\omega$, the random walk $(Z_{n}(\omega)o)$ generated by $\mu$ is $\lambda$-regular. In particular, $(Z_{n}o)$ converges almost surely to a regular point of the visual boundary contained in the chamber at infinity $\psi(\text{bnd}(\omega))$, where $\psi$ is the boundary map given by Theorem 1.1, and almost surely,
$\lim_{n}\frac{1}{n}\theta(o,Z_{n}o)=\lambda.$
We point out that the moment assumption in the previous theorem is not optimal. In a subsequent work, we will show that without moment condition, $(Z_{n}(\omega)o)$ converges almost surely to the chamber at infinity $\psi(\text{bnd}(\omega))$ in the polyhedral sense, that is, for the topology of the polyhedral compactification (the analogue of the maximal Satake compactification of symmetric spaces), see [Rou23, §3.4]. We conjecture that the Theorem 1.4 holds if $\mu$ is only assumed to have finite first moment
[…]
and that under a finite second moment assumption, one can prove a central limit theorem using the strategy of Benoist and Quint in [BQ16b]. Let us also mention that this result relies on the specific geometry of $\tilde{A}_{2}$-buildings and the fact that we used a symmetric measure: Theorem 1.4 no longer holds for $\tilde{C}_{2}$ or $\tilde{G}_{2}$-buildings even in the algebraic case, as there exist Zariski-dense discrete subgroups of algebraic groups over non-Archimedean local fields whose limit cones lie in proper facets of the model Weyl chamber [Qui02].
The aforementioned results can be translated in the algebraic setting. For a general presentation to Bruhat-Tits buildings, we refer to [BT84], [Wei09] and [KP23]. Let $\mathbf{G}$ be a connected semisimple (or even reductive quasi-split) algebraic group defined over a separable non-Archimedean field $K$ or, more generally, a division algebra. Consider the associated Bruhat-Tits building $X=\Delta^{\text{BT}}(\mathbf{G})$, as defined in [BT84]. If the valuation is discrete, then $X$ is automatically complete, but some care is needed when the valuation is dense. Precise conditions for $X$ to be metrically complete in this case are given in [MSSS13]. Namely, the field or division algebra must be spherically complete. In any case, if the field is just assumed complete, then the spherical building at infinity $X^{\infty}=\Delta(\mathbf{G})$ of $X$ is well-defined, and we have the identification
$\text{ch}(\Delta(\mathbf{G}))\simeq\mathbf{G}(K)/\mathbf{P}(K),$
where $\mathbf{P}$ is a minimal parabolic subgroup of $\mathbf{G}$.
Under these assumptions, our results translate as the following:
Theorem 1.5.
Let $\mathbf{G}$ be an absolutely almost simple algebraic group over a separable and spherically complete field with valuation $v:K^{\times}\to\mathbb{R}$. We assume that the Weyl group of $\mathbf{G}$ is of type $A_{2}$ (for instance, $\mathbf{G}=\text{SL}_{3}(K)$ or $\text{PGL}_{3}(K)$). Let $\Gamma$ be a locally compact second countable group and let $\Gamma\to\mathbf{G}$ be a group representation such that the induced $\Gamma$-action on $\Delta^{BT}(\mathbf{G})$ is non-elementary (for instance, $\pi(\Gamma)$ is Zariski-dense and unbounded), and let $(B,\nu)$ be a $\Gamma$-boundary. Then there exists a unique measurable $G$-map
$B\rightarrow\mathbf{G}(K)/\mathbf{P}(K),$
for $\mathbf{P}$ a minimal parabolic subgroup of $\mathbf{G}$.
In particular, if $\Gamma$ is a lattice in a semisimple $\mathbb{K}$-algebraic group $\mathbf{H}$ for which $\mathbf{P_{0}}$ is a minimal parabolic subgroup, and if $\pi:\Gamma\to\mathbf{G}$ is a group homomorphism with Zariski-dense unbounded image, then $\mathbf{H}(\mathbb{K})/\mathbf{P_{0}}(\mathbb{K})$ is a $\Gamma$-boundary and there exists a unique $\Gamma$-equivariant measurable map
$\mathbf{H}(\mathbb{K})/\mathbf{P_{0}}(\mathbb{K})\rightarrow\mathbf{G}(K)/\mathbf{P}(K).$
This result for local fields is essentially due to Zimmer, and it is the first step in his approach of the proof of Margulis super-rigidity [Zim84, Chapter 5].
We can also translate the statement about the convergence of the random walk in Theorem 1.4. We denote by $\mathfrak{a}^{+}$ a Weyl chamber associated to a choice of positive roots, and by
$\kappa:\mathbf{G}\to\mathfrak{a}^{+}$
the Cartan projection given by the Cartan decomposition associated to a choice of a maximal compact subgroup. For more details on these notions, see for example [BQ16a, Chapter 5].
Theorem 1.6.
Consider the same assumptions as in Theorem 1.5. If $\mu$ is an admissible measure on $\Gamma$, then there exists a unique $\mu$-stationary probability measure on the Furstenberg boundary $\mathbf{G}(K)/\mathbf{P}(K)$. Moreover, the Lyapunov spectrum of the random walk $(Z_{n}(\omega))$ is simple: there exists $\lambda\in\mathfrak{a}^{++}$ such that almost surely,
$\lim_{n}\frac{1}{n}\kappa(Z_{n}(\omega))=\lambda.$
It should be noted that in the non-Archimedean case, the classification of stationary measures on the flag variety $\mathbf{G}(K)/\mathbf{P}(K)$ is more subtle than in the real case. For instance, for $K=\mathbb{Q}_{p}$, $p$ a prime number, there exist measures $\mu$ such that the semigroup $\Gamma_{\mu}$ generated by $\mu$ is Zariski-dense, and for which there exist more than one $\mu$-stationary measures on the flag variety. It is the case for instance if $\Gamma_{\mu}$ is a small open compact subgroup of $\mathbf{G}(K)$, see [BQ14, Remark 4.18]. Here however, the representation is also unbounded, and the uniqueness points to the fact that the action is actually proximal, and this fact will get transparent in the paper [LBLS24].
1.3. Reduction when the building is not complete separable
We end this introduction by discussing the case where the building $X$ is no longer assumed separable and complete. In the paper [LBLS24], we prove that in this situation, the aforementioned results still hold, provided the group $G$ is discrete countable and acts in a non-elementary way. More precisely, we prove that if the $\tilde{A}_{2}$-building $X$ is not assumed separable, there exists $Y\subseteq X$ a sub-building of type $\tilde{A}_{2}$, which is separable and $G$-invariant. The $G$-action on $Y$ is clearly non-elementary. We can then, using ultralimits of metric spaces and results of Kleiner and Leeb [KL97] embed $Y$ in a metrically complete $\tilde{A}_{2}$-building $\hat{Y}$. The metric completion $\bar{Y}$ naturally sits in this completion: it is separable, metrically complete and endowed with a natural $G$-action by isometries. In [LBLS24], we prove that the visual boundary of $\bar{Y}$ is actually $\partial_{\infty}Y$, therefore there the $G$-action on $\bar{Y}$ remains non-elementary. On the other hand $\bar{Y}$ is not necessarily a building. But a careful look at the proofs here shows that all the results remain true in this setting. Specifically, consider the assumptions:
  • $X$ is a separable metrically complete CAT(0) space, such that $\partial_{\infty}X$ can be given the structure of a spherical building of type $A_{2}$, and $X$ is contained in a (possibly non-separable) metrically complete $\tilde{A}_{2}$-building $\hat{X}$.
  • The group $G$ is discrete countable, and acts by isometries on $X$ without bounded orbit in $X$ (equivalently, without fixed point) nor finite orbit in $\partial_{\infty}X$.
As explained before, this situation applies when a discrete countable group $G$ acts on any $\tilde{A}_{2}$-building $X^{\prime}$, with $X$ being constructed by the reduction arguments above. Then the following result is true:
Theorem 1.7.
Under the above assumptions, the results stated in Theorems 1.1, 1.2, 1.4 and Proposition 1.3 hold.
In particular, for the algebraic case, it means that our results remain true if the field is only assumed to be complete, and no longer only for fields that are spherically complete and separable.
1.4. Structure of the proof
An essential ingredient for the existence of the boundary map $B\rightarrow\text{ch}(X^{\infty})$ is a result due to Bader, Duchesne and Lécureux in [BDL16]. The authors prove that for every non-elementary $G$-action on a CAT(0) space of finite telescopic dimension, there exists a boundary map $\psi:B\rightarrow\partial_{\infty}X$, for $B$ a $G$-boundary. We investigate the characteristics of this map when $X$ is a building of type $\tilde{A}_{2}$, using the nice ergodic properties of $B$ and the geometric structure of $\partial_{\infty}X$. If almost surely, $\psi(b)$ is a regular point of the boundary, i.e. is not represented by a panel, we get a $G$-map $B\to\text{ch}(X^{\infty})$, which is what we want. This leads us to study the case where the target space is the set of vertices at infinity. We therefore assume that there is a measurable $G$-map $B\to X^{\infty}_{1}$ to the vertices at infinity of type 1. To every vertex $u$ of the spherical building at infinity, one can associate a tree $T_{u}$, called the panel tree at $u$, and whose ends $\partial T_{u}$ correspond canonically to the chambers at infinity in the residue $\text{Res}(u)$. We will consider the set the fibered space $T_{1}=\sqcup_{u\in X^{\infty}_{1}}T_{u}$ of panel trees over the vertices of type 1, and show that they can be given a Borel structure. In this case, a theorem due to Duchesne [Duc13], generalizing Adams-Ballmann equivariant theorem [AB98] gives exactly three possibilities: either there is $G$-map $B\to\partial T_{1}$ with values in the boundary field of $T_{1}$, or a $G$-map $B\to T_{1}$, or a $G$-map $B\to\partial_{2}T_{1}$ with values in the geodesic lines of the field $T_{1}$. Using the strong ergodic properties of $G$-boundaries and the geometric features of $\tilde{A}_{2}$-buildings, we rule out the latter two possibilities, and conclude with the bijection between the ends of the panel trees and the residues of vertices a infinity.
Once we know that there is a unique $\mu$-stationary measure $\nu$ on the boundary, we can proceed to study the random walk $(Z_{n}o)$. We use the notion of $\lambda$-ray developed by V. Kaimanovich [Ki87] for symmetric spaces and J. Parkinson and W. Woess in the context of buildings [PW15]. We then apply a combination of results by H. Izeki [Ize23] and A. Karlsson and G. Margulis [KM99] to our context. We conclude with the particular behavior of the opposition involution in $\tilde{A}_{2}$-buildings.
Structure of the article
In Section 2, we review the definition of (non-discrete) affine building as well as some related objects, including the panel trees associated to vertices at infinity. Section 3 presents some notions of boundary theory such as measurable metric fields and relative metric ergodicity, and states some key results in this field. In Section 4, we ensure that there exist nice measurable structures for some fibered spaces, in order to apply the results from Section 3. We prove Theorem 1.1 in Section 5, and in Section 6, we prove the main Theorem 1.2, and we derive Proposition 1.3. In Section 7, we show that the random walk associated to an admissible measure with finite second moment on $G$ converges almost surely a regular point of the boundary.
Acknowledgements.
The author is very grateful to J. Lécureux, who originated many of the ideas implemented here and for helpful discussions.
2. Affine buildings of type $\tilde{A}_{2}$
In this section, we define non-discrete affine buildings, and we introduce the objects that we will use throughout the paper.
2.1. Non-discrete affine buildings
Let $V$ be a finite dimensional Euclidean vector space, and let $W_{0}$ be a finite reflection group acting on $V$, that is, a finite group of linear isometries of $V$ generated by (linear) reflections. The fixed point set of a reflection in $W_{0}$ is called a linear wall of $V$. There are finitely many walls in $V$. A vector Weyl chamber of $(W_{0},V)$ is a connected component of the complement in $V$ of the union of all the $W_{0}$-walls. The closure of a vector Weyl chamber is a closed vector Weyl chamber. Any such vector Weyl chamber is a polyhedral cone, that is, the intersection of a finite number of half-spaces. A vector facet (resp. closed vector facet) of $(W_{0},V)$ is a facet (resp. closed facets) of some vector Weyl chamber: it is the intersection of a Weyl chamber with a finite number of walls. A vector facet of codimension 1 is called a vector panel. As $W_{0}$ acts on the set of vector Weyl chambers transitively, we choose $\mathfrak{a}^{+}$ to be some closed vector Weyl chamber, so that $\mathfrak{a}^{+}$ is a fundamental domain for the $W_{0}$-action on $V$. We call $\mathfrak{a}^{+}$ the fundamental (closed) Weyl chamber.
Now let $W$ be a group of affine isometries of $V$ generated by affine reflections, and assume that the group $W_{0}$ generated by the vector parts of the elements of $W$ is finite. We call $W$ an affine reflection group, and $(W,V)$ an affine reflection system. Any such group can be written $W=W_{0}\ltimes T$, where $W_{0}$ is a finite reflection group and $T$ is a translation group on $V$. A sector (or affine Weyl chamber) is then defined to be the translate of a vector Weyl chamber by an element of $T$. Similarly, we shall call affine walls, affine facet, affine panels etc. the translates of the vector walls, vector facets, vector panels etc. We say that the affine reflection group $W$ is discrete if $T$ is discrete in $V$. Otherwise, we say that the reflection group is non-discrete.
We can now turn to the definition of an affine building.
Definition 2.1.
Let $(W,V)$ be an affine reflection system. Let $X$ be a set, and let $\mathcal{A}$ be a collection of injective charts of $V$ into $X$, which we call an atlas. Each such injection is called a chart, or marked apartment, and the image $A$ of $V$ by an injection is called an apartment. We say that $(X,\mathcal{A})$ is a affine building modelled after $(W,V)$ if the following axioms are verified.
  1. (A1)
    The atlas $\mathcal{A}$ is invariant by pre-composition with $W$.
  2. (A2)
    Given two charts $f,f^{\prime}:V\to X$ with $f(V)\cap f^{\prime}(V)\neq\emptyset$, then $U:=f^{-1}(f^{\prime}(V))$ is a closed convex subset of $V$, and there exists $w\in W$ such that $f|_{U}=f^{\prime}\circ w|_{U}$.
  3. (A3)
    For any pair of points $x,y\in X$, there is an apartment containing both.
Axioms (A1)-(A3) imply the existence of a well-defined distance function $d:X\times X\to\mathbb{R}_{+}$, such that the distance between any two points is the $d_{V}$-distance between their pre-image under any chart containing both. A Weyl chamber (or sector) in $X$ is the image of an affine Weyl chamber under some chart $f\in\mathcal{A}$.
  1. (A4)
    Given two Weyl chambers $S_{1},S_{2}$ in $X$, there exist sub-Weyl chambers $S_{1}^{\prime}\subseteq S_{1},S_{2}^{\prime}\subseteq S_{2}$ such that $S_{1}^{\prime}$ and $S_{1}^{\prime}$ are contained in the same apartment.
  2. (A5)
    For any apartment $A$ and $x\in X$, there exists a retraction $\rho_{A,x}:X\to A$ such that $\rho_{A,x}$ does not increase distance and $\rho^{-1}_{A,x}(x)=\{x\}$.
If the previous axioms are verified, then the distance function $d$ is a metric and the building $(X,d)$ is a metric space [Par00, Proposition 1.3]. In fact, this metric is CAT(0).
If $W$ is discrete and $W_{0}$ is irreducible, that is, can not be written as a non-trivial direct product $W_{1}\times W_{2}$ acting on the vector space $V=V_{1}\oplus V_{2}$, then the building $(X,\mathcal{A})$ modelled after $(W,V)$ is actually the geometric realization of a simplicial affine building. If the affine reflection group is not discrete, we say that the building $(X,\mathcal{A})$ modelled after $(W,V)$ is non-discrete.
Let $(X,\mathcal{A})$ be an affine building modelled after $(W,V)$, let $\mathfrak{a}^{+}$ be a fundamental Weyl chamber of $(W,V)$, and let $x,y\in X$. By the axioms (A1)-(A3), there exists a marked apartment $f$ sending the fundamental closed Weyl chamber $\mathfrak{a}^{+}$ to a Weyl chamber in $X$ based at $x$ and containing $y$. The type $\theta(x,y)$ of the Euclidean segment $[x,y]$ is then unique vector in $\mathfrak{a}^{+}$ such that $y=f(\theta(x,y))$.
A facet (resp. closed facet) of $X$ is the image of an affine facet (resp. of an affine closed facet) of $(W,V)$ under a chart $f\in\mathcal{A}$. Recall that an affine facet of $V$ is written $v+F_{0}$, for $F_{0}$ a vector facet in $V$. We then say that the facet $F$ of $X$ is based at $x$ if there exists $f\in\mathcal{A}$ such that $F=f(v+F_{0})$ and $x=f(v)$. Up to pre-composition by an element of $W$, we can assume that $F_{0}$ is a facet of $\mathfrak{a}^{+}$, and we say that the type $\theta(F)$ of the facet $F$ is $F_{0}$.
We say that a point $x\in X$ is special if it is the image of a special point $v$ in $(W,V)$, i.e. a point for which the stabilizer $W_{v}:=\text{Stab}_{W}(v)$ is conjugated to $W_{0}$. We denote by $X_{s}$ the set of special points in $X$. If $T=V$, which the reader can assume for convenience, every point in an $\tilde{A}_{2}$-building is special. Also, for convenience we will only consider case where the system of apartments $\mathcal{A}$ is maximal, i.e. is not contained properly in a system of apartment for $X$.
2.2. Visual boundary and spherical building at infinity
Let $(X,\mathcal{A},d)$ be an affine building modeled after $(W,V)$, and let $\mathfrak{a}^{+}$ be a fundamental Weyl chamber of $(W,V)$. As a CAT(0) space, $X$ has a natural bordification given by equivalence class of rays, two rays being equivalent if they are at finite Hausdorff $d$-distance. We call this bordification the visual or conical bordification, and denote it by $\overline{X}=X\cup\partial_{\infty}X$. When $X$ is separable, the visual boundary $\partial_{\infty}X$ is endowed with a natural metrizable topology and isometries of $X$ extend to homeomorphisms on the boundary. As $X$ is an affine building, we can refine this structure in the following way. We say that two facets $F$ and $F^{\prime}$ are parallel if they are at finite Hausdorff distance. This is an equivalence relation, and a class of facets is called a facet at infinity. The facet at infinity determined by a facet $F$ is denoted by $F^{\infty}$. Equivalence classes of facets of dimension 1, of codimension 1 and of maximal dimension are called vertices, panels and chambers at infinity respectively. We denote the set of chambers at infinity by $\text{ch}(X^{\infty})$. Let $F^{\infty}_{1}$ and $F^{\infty}_{2}$ be two facets at infinity. We say that $F^{\infty}_{1}$ dominates $F^{\infty}_{2}$ if for any facet $F_{1}\in F^{\infty}_{1},F_{2}\in F^{\infty}_{2}$, we have
$\sup_{x\in F_{2}}d(x,F_{1})<\infty.$
The set $X^{\infty}$ of all facets at infinity endowed with the domination relation is a spherical building, whose apartments are in one-to-one bijection with the apartments of $X$, see [Par00, Propriété 1.7]. Note that for $x$ a special vertex, and $F^{\infty}$ a facet at infinity, there is an affine facet based at $x$ in the equivalence class of $F^{\infty}$, which we denote by $Q(x,F^{\infty})$.
Let $x,y\in X$ be special points of the building $X$. By the buildings axioms, there is a marked apartment $f\in\mathcal{A}$ containing both, and again by the building axioms, we can assume that $f(0)=x$ and $f^{-1}(y)\in\mathfrak{a}^{+}$. We say that the segment $[x,y]$ is regular if the type $\theta(x,y)$ is regular, i.e. $\theta(x,y)=f^{-1}(y)\in\mathfrak{a}^{++}$. If $[x,y]$ is regular, then both $\theta(x,y)$ and $\theta(y,x)$ are regular. Notice that in this case, $f(\mathfrak{a}^{+})$ is a Weyl chamber of the building $X$, contained in $A$ and containing $y$. Denote by $C_{y}$ the chamber at infinity that it represents. Similarly, $f(\theta(x,y)-\mathfrak{a}^{+})$ is a Weyl chamber containing $x$ representing a chamber $C_{x}\in\text{ch}(X^{\infty})$. They are opposite, that is, there exists a unique apartment joining them, namely $A$.
Denote by $w_{0}$ the long element of the finite Weyl group $W_{0}$ associated with the affine reflection group $W$. For $\lambda\in\mathfrak{a}$, the opposition involution $\iota:\mathfrak{a}^{+}\to\mathfrak{a}^{+}$ is defined by
$\iota(\lambda)=w_{0}(-\lambda).$
A detailed account on how to give a topology on the set $X\cup\text{ch}(X^{\infty})$, and even on $X\cup X^{\infty}_{\tau}$, where $X^{\infty}_{\tau}$ represents the set of simplices of type $\tau$ of the spherical building at infinity, is given in [Rou23]. We will not give much details on the topology, but we summarize here a few results.
The set of chambers at infinity $\text{ch}(X^{\infty})$ can be endowed with a topology, which makes it a totally disconnected space. A basis of open neighborhoods in $\text{ch}(X^{\infty})$ is given by
(1) $U_{x}(y):=\{C\in\text{ch}(X^{\infty})\,|\,y\in Q(x,C)\}\subseteq\text{ch}(X^{\infty}),$
for $x,y\in X$.
The following proposition summarizes some properties of this bordification.
Proposition 2.2 ([Rou23, §3.2]).
Let $X$ be any Euclidean building. Then there is a topology on $X\cup\text{ch}(X^{\infty})$ for which a basis of open sets of the chambers at infinity is given by the sets (1). It agrees with the CAT(0) topology on $X$. This topology is first-countable and Hausdorff. For every boundary point $\xi\in\partial_{\infty}X$ that is strictly supported on a chamber, one can associate canonically a chamber $C_{\xi}\in\text{ch}(X^{\infty})$. This map is a homeomorphism with respect to the restriction of the visual topology on $\partial_{\infty}X$ to such points.
From the geometric viewpoint, the boundary $\partial_{\infty}X$ of a CAT(0) space $X$ can be endowed with the Tits metric, with which it becomes a $\text{CAT}(1)$ space and on which $\operatorname{Isom}(X)$ still acts by isometries. The matter of existence of fixed points for a given group action on a $\text{CAT}(1)$ space has been widely studied, see [BL05], [CL10] and [CM13]. The following is taken from [BL05], and will be useful later.
Proposition 2.3 ([BL05, Proposition 1.4]).
Let $X$ be a CAT(1) space of finite dimension and of radius $r\leq\pi/2$. Then $X$ has a circumcenter which is fixed by every isometry of $X$.
2.3. Residue building based at a vertex
In this section, we show that even in non-discrete buildings, there is a good notion of (local) alcoves and faces.
Definition 2.4.
Let $F$ and $F^{\prime}$ be two facets based at a special vertex $o\in X$. We say that $F$ and $F^{\prime}$ have the same germ at $o$ if their intersection is an open neighborhood of $o$ in both $F$ and $F^{\prime}$.
This is an equivalence relation, and we denote the class of a given facet $F$ by $\text{germ}_{o}(F)$. We say that $\text{germ}_{o}(F)$ dominates $\text{germ}_{o}(F^{\prime})$ if the closed facet $\overline{F}$ contains an open neighborhood of $o$ in $F^{\prime}$. The set of all germs at $o$ that are not reduced to $\{o\}$, equipped with the partial order given by this domination relation is denoted $\Sigma_{o}X$. It is a simplicial spherical building [Par00, Corollary 1.11]
Definition 2.5.
The simplicial complex $\Sigma_{o}X$ is called the residue building or the building of germs based at $o$.
Simplices of maximal dimension of $\Sigma_{o}X$ are called (local) alcoves. Their codimension-1 faces are called local panels.
For every special vertex $x\in X$, we have a canonical morphism of simplicial complexes
(2) $\Sigma_{x}:X^{\infty}\to\Sigma_{x}X$
sending any facet at infinity $F^{\infty}$ to $\text{germ}_{x}(Q(x,F^{\infty}))$.
2.4. Retractions and automorphisms
Let $X$ be an affine building, $A$ an apartment, $x\in A$ a vertex and let $S$ be a Weyl chamber in the apartment $A$, based at $x$. Then there exists a unique retraction map $\rho_{A,S}:X\to A$ such that $\rho_{A,S}$ preserves the distance on any apartment containing a sub-Weyl chamber of $S$, see [Par00, Proposition 1.20]. Moreover, $\rho_{A,S}$ does not increase distances. We call $\rho_{A,S}$ the (sector-based) retraction of $X$ on $A$ based at $S$.
In this paper, we will only consider automorphisms of buildings in the sense of Rousseau [Rou23, §2.1.13], that is, isometries that preserve the structure of walls. Notice that in our case, we shall only consider irreducible buildings with the maximal system of apartments, therefore isometries of $X$ are often automatically automorphisms: this is the case in particular if the building is thick, see [Par00, Proposition 2.27] (in Parreau’s paper, the notion of automorphisms is more restrictive, it is automatically type-preserving). The group of automorphisms $\text{Aut}(X)$ contains a finite-index subgroup formed by automorphisms that are type-preserving, i.e. automorphisms $f$ such that for any segment $[x,y]$, $\theta(x,y)=\theta(f(x),f(y))$, see [Rou23, §2.4.6.1]. Hereafter, we always implicitly make the assumption that the $G$-action is type-preserving.
Moreover, any type-preserving automorphism induces a type-preserving automorphism of the spherical building at infinity $X^{\infty}$. With the topology on $\text{ch}(X^{\infty})$, an automorphism of $X$ can be extended to a homeomorphism on $\text{ch}(X^{\infty})$.
2.5. $\tilde{A}_{2}$-buildings
Let now $V$ be a Euclidean vector space of dimension 2, and let $W=W_{0}\ltimes T$ be an irreducible affine reflection group acting on $V$, with $W_{0}$ an irreducible spherical reflection group and $T$ acting as a non-trivial translation group. We assume that $W$ is essential, that is, $W_{0}$ has no non-trivial fixed point in $V$. Let $X$ an affine building of modeled after $(W,V)$. We say that $X$ is of type $\tilde{A}_{2}$ if $W_{0}$ is a spherical Coxeter group of type $A_{2}$. Let $\mathbb{A}$ be a fundamental apartment of $X$. In the discrete case, $\mathbb{A}$ is the 2-dimensional simplicial complex afforded by the tessellation of the Euclidean plane $V$ by equilateral triangles, and any apartment of $X$ is a subcomplex isomorphic to $\mathbb{A}$.
In the $\tilde{A}_{2}$-case, the spherical building at infinity $X^{\infty}$ is a graph whose vertices are the vertices at infinity, and for which there is an edge between two vertices if they are dominated by a common chamber at infinity. The set of edges in $X^{\infty}$ is still denoted by $\text{ch}(X^{\infty})$: it is the set of equivalence classes of sectors. In the same way, $X^{\infty}$ is endowed with a coloring that divides its vertices into two types, and the corresponding subsets are denoted $X^{\infty}_{1}$ and $X^{\infty}_{2}$.
2.6. Panel trees
Recall that a boundary point $v\in\partial_{\infty}X$ is an equivalence class of rays, two rays $r_{1}$ and $r_{2}$ being equivalent, for which we write $r_{1}\sim r_{2}$, if they contain subrays that lie in a common apartment and are parallel in this apartment. We will say that two geodesic rays $r_{1}$ and $r_{2}$ are strongly asymptotic, and write $r_{1}\simeq r_{2}$, if their intersection contains a geodesic ray. For two equivalent geodesic rays $r_{1}$ and $r_{2}$ that represent the boundary point $v\in\partial_{\infty}X$, we define their distance to be:
[…]
Note that it defines a distance, and that two strongly asymptotic rays $r_{1}$ and $r_{2}$ satisfy $d_{v}(r_{1},r_{2})=0$. Then this distance does not depend on the $\simeq$-strongly asymptotic classes of ray among rays from the same $\sim$-equivalence class. The following proposition describes the panel-trees when the affine building is simplicial.
Proposition 2.6 ([BCL19, Proposition 2.13]).
Let $X$ be a simplicial thick $\tilde{A}_{2}$-building. The metric space $(T_{v},d_{v})$ of asymptotic classes of rays in the class of a vertex at infinity $v$ is a thick tree, and if $X$ is locally finite of order $q$, $T_{v}$ is regular of order $q+1$.
For an affine building of dimension 2 and $v$ a vertex at infinity, the space $T_{v}$ is a thick $\mathbb{R}$-tree (an affine $\tilde{A}_{1}$-building) called the panel tree at $v$. Note that this notion can be generalized in higher dimension: in this case, we should replace vertices at infinity with panels, see for instance [Wei09]. Actually, to any facet at infinity, one can attach an affine building structure, see [RT21, Section 4.3] and [CL11, Section 5]. For $v$ a vertex at infinity, we denote by $\text{Aut}(X)_{v}$ the group of automorphisms of $X$ that stabilize $v$.
Note that for all vertex at infinity $v\in X^{\infty}$, there is a well-defined and continuous application defined by
$\pi_{v}:X$$\longrightarrow$ $T_{v}$
[…]
where $[Q(x,v)]$ is the class (for the strongly asymptotic relation) of the geodesic ray based at $x$, in the direction of $v$.
A key property, which we state in Proposition 2.7 below, is that the space of ends of a given panel tree $T_{v}$ is in bijection with the residue $\text{Res}(v)$ of the chambers at infinity containing $v$. This fact was known to Tits [Tit86], and was used in its classification of affine buildings. Let us describe this application for any affine building. Let $F^{\infty}$ be a panel at infinity, and let $S$ be a sector in $X$, one of whose faces represents $F^{\infty}$. In other words, the parallel class $[S]_{\sim}$ of $S$ belongs to the residue $\text{Res}(F^{\infty})$ of $F^{\infty}$ in the building at infinity $X^{\infty}$. We have a well-defined homomorphism of affine buildings
(4) $\pi_{F^{\infty}}:X$$\longrightarrow$ $T_{F^{\infty}}$
[…]
where $[Q(x,F^{\infty})]$ is the class of the panel based at $x$, in the direction of $F^{\infty}$. In fact, this is valid for any facet at infinity (not only of codimension 1) when we replace the panel tree $T_{F^{\infty}}$ by the corresponding affine building at infinity $X(F^{\infty})$, see [RT21, Section 4.3].
Let $S,S^{\prime}$ be Weyl chamber in $X$ such that $F^{\infty}\in\partial_{\infty}S$ and $F^{\infty}\in\partial_{\infty}S^{\prime}$. In other words, the parallel classes of $S$ and $S^{\prime}$ are in the residue of $F^{\infty}$. If we assume furthermore that $S$ and $S^{\prime}$ are parallel, then there is a sub-sector $S^{\prime\prime}\subseteq S\cap S^{\prime}$, and $\pi_{F^{\infty}}(S^{\prime\prime})$ is a common sub-sector of the sectors $\pi_{F^{\infty}}(S)$ and $\pi_{F^{\infty}}(S^{\prime})$ in the affine building $X(F^{\infty})$. This allows to define a surjective map
$\phi_{F^{\infty}}$ ::: $\text{Res}(F^{\infty})\longrightarrow\partial T_{F^{\infty}}$
$[S]_{\sim}\longmapsto[\pi_{F^{\infty}}(S)]_{\sim},$
where $\partial T_{F^{\infty}}$ is to be understood as the set of chambers at infinity of the affine $\tilde{A}_{1}$-building $T_{F^{\infty}}$. It was already known to Tits [Tit86, Proposition 4] that this map is a bijection. We now prove that $\phi_{F^{\infty}}$ is a homeomorphism, when we endow $\text{Res}{(F^{\infty})}$ and $\partial T_{F^{\infty}}$ with the natural topology for chambers at infinity on affine buildings. This result is not surprising for experts, but we did not find any reference in the literature in the non-discrete case. In the discrete case, it is proven in [RT21, Lemma 4.2].
Recall that if $X$ is a building, a topology on the set of chambers at infinity $\text{ch}(X^{\infty})$ is given by the basis of open sets: for all $x,y\in X$ special vertices of $X$,
$U_{x}(y):=\{C\in\text{ch}(X^{\infty})\,:\,y\in Q(x,C)\},$
where $Q(x,C)$ is the unique sector of $X$ based at $x$, in the parallel class of $C$.
Proposition 2.7.
Let $F^{\infty}$ be a panel in the spherical building at infinity $X^{\infty}$. Then the map
$\phi_{F^{\infty}}:\text{Res}(F^{\infty})\longrightarrow\partial T_{F^{\infty}}$
is a $\text{Aut}(X)_{v}$-equivariant homeomorphism between the residue $\text{Res}(F^{\infty})$ and the space of ends of the panel tree $T_{F^{\infty}}$.
Proof.
The fact that $\phi_{F^{\infty}}$ is a bijection is well-known, see for example [Tit86, Proposition 4] or [Wei09, Proposition 11.22]. It is $\text{Aut}(X)_{F^{\infty}}$-equivariant by construction.
Let $\xi,\eta$ be two points of $T_{F^{\infty}}\cap\pi_{F^{\infty}}(X_{s})$, and let $\mathcal{C}\in\partial T_{F^{\infty}}$. In order to avoid confusion, we denote by $Q_{F^{\infty}}(\xi,\mathcal{C})$ the sector (which is a ray since $T_{F^{\infty}}$ is a tree) based at $\xi$ in the direction of the end $\mathcal{C}$. We shall write $U^{\infty}_{\xi}(\eta)$ to denote the set
$\{\mathcal{C}\in\partial T_{F^{\infty}}\,:\,\eta\in Q_{F^{\infty}}(\xi,\mathcal{C})\}.$
Let $C\in\phi_{F^{\infty}}^{-1}(U^{\infty}_{\xi}(\eta))$ be a chamber at infinity. The sector $Q_{F^{\infty}}(\xi,\phi_{F^{\infty}}(C))$ belongs to an apartment $A_{\infty}$ of the panel-tree $T_{F^{\infty}}$. Such an apartment can be written $\pi_{F^{\infty}}(A)$, for $A$ an apartment in $X$ such that $\partial_{\infty}A$ contains $F^{\infty}$. Let $x\in X_{s}\cap A$ be such that $\pi_{F^{\infty}}(x)=\xi$. By surjectivity, there exists a sector $S$ in $A$ based at $x$ such that $\pi_{F^{\infty}}(S)$ is a sector of $T_{F^{\infty}}$ representing $\phi_{F^{\infty}}(C)$. Let $y$ be a special vertex in $S$ be such that $\pi_{F^{\infty}}(y)=\eta$. Then by injectivity,
$Q_{F^{\infty}}(\xi,\phi_{F^{\infty}}(C))=\pi_{F^{\infty}}(S).$
Now for any $C^{\prime}\in\text{Res}(F^{\infty})\cap U_{x}(y)$, we have $y\in Q(x,C^{\prime})$ and thus $\eta\in Q_{F^{\infty}}(\xi,\phi_{F^{\infty}}(C))$. In other words, $U_{x}(y)\cap\text{Res}(F^{\infty})$ is contained in $\phi_{F^{\infty}}^{-1}(U^{\infty}_{\xi}(\eta))$, and $\phi_{F^{\infty}}$ is continuous.
Conversely, let $x,y\in X_{s}$ be special vertices and let $\xi\in\phi_{F^{\infty}}(U_{x}(y)\cap\text{Res}(F^{\infty}))$. Take $C\in U_{x}(y)\cap\text{Res}(F^{\infty})$ such that $\phi_{F^{\infty}}(C)=\mathcal{C}$. Define $\xi=\pi_{F^{\infty}}(x)$ and $\eta=\pi_{F^{\infty}}(y)$. By definition, $Q_{F^{\infty}}(\xi,\mathcal{C})$ is a sector of $T_{F^{\infty}}$. But by injectivity, $Q_{F^{\infty}}(\xi,\mathcal{C})=\pi_{F^{\infty}}(Q(x,C))$ and $y\in Q(x,C)$. Therefore, $\eta\in Q_{F^{\infty}}(\xi,\mathcal{C})$ and $\mathcal{C}\in U^{\infty}_{\xi}(\eta)$. ∎
3. Metric fields and boundary theory
3.1. Boundary theory
Let $G$ be a locally compact second countable group, and let $\mu$ be a probability measure on $G$. Let $(\Omega,\mathbb{P})$ be the probability space $(G^{\mathbb{N}},\mu^{\otimes\mathbb{N}})$, with the product $\sigma$-algebra. The space $\Omega$ is called the space of forward increments. The application
$(\omega=(\omega_{i})_{i\in\mathbb{N}},n)\in\Omega\times\mathbb{N}\mapsto Z_{n}(\omega)=\omega_{1}\omega_{2}\dots\omega_{n},$
defines the random walk on $G$ generated by the measure $\mu$.
Let $Y$ be a topological space, endowed with the $\sigma$-algebra of Borel sets. Let $G\curvearrowright Y$ be a continuous action on $X$, and let $\nu\in\text{Prob}(Y)$ be a probability measure on $Y$. We define the convolution probability measure $\mu\ast\nu$ as the image of $\mu\times\nu$ under the action $G\times Y\rightarrow Y$. In other words, for $f$ a bounded measurable function on $Y$,
[…]
We will write $\mu_{m}=\mu^{\ast m}$ the $m$-th convolution power of $\mu$, where $G$ acts on itself by left translation $(g,h)\mapsto gh$.
Definition 3.1.
We say that $\mu\in\text{Prob}(G)$ is admissible if its support $\operatorname{\text{supp}}(\mu)$ generates $G$ as a semigroup, and there exists $n\in\mathbb{N}$ such that $\mu_{n}$ and the Haar measure on $G$ are non-singular.
From now on, we will always assume that the probability measure on $G$ is admissible. Let us turn to the main notion of this section.
Definition 3.2.
A probability measure $\nu\in\text{Prob}(Y)$ is $\mu$-stationary if $\mu\ast\nu=\nu$. In this case, we say that $(Y,\nu)$ is a $(G,\mu)$-space.
Notice that if $Y$ is a Polish space, the set of probability measures $\text{Prob}(Y)$ can be endowed with the metrizable weak-$\ast$ topology. The following statement, due to H. Furstenberg, is one of the fundamental theorems of boundary theory.
Theorem 3.3 ([Fur73, Lemma 1.33]).
Let $G$ be a locally compact second countable group, $\mu\in\text{Prob}(G)$ and admissble measure and $(Z_{n}(\omega))$ be the random walk on $G$ associated to the measure $\mu$. Let $Y$ be a metrizable space on which $G$ acts by isometries, and let $\nu$ be a $\mu$-stationary measure on $Y$. Then, there exists an essentially well-defined $G$-equivariant measurable map $\omega\mapsto\nu_{\omega}\in\text{Prob}(Y)$ such that $\mathbb{P}$-almost surely, $Z_{n}(\omega)\nu\rightarrow\nu_{\omega}$ in the weak-$\ast$ topology. Moreover, we have the decomposition
[…]
Definition 3.4.
A $(G,\mu)$-space $(Y,\nu)$ is a $(G,\mu)$-boundary if for $\mathbb{P}$-almost every $\omega\in\Omega$, $\nu_{\omega}$ is a Dirac measure.
The study of $(G,\mu)$-boundaries has strong connections with the existence of harmonic functions and Poisson transforms on a group, but these results will be omitted here. We refer to [Fur02] for more information on these subjects.
For $(Y,\nu)$ and $(Y^{\prime},\nu^{\prime})$ non-singular $(G,\mu)$ spaces, we say that a measurable map $\pi:(Y,\nu)\to(Y,\nu)$ is a factor map if $\pi_{\ast}\nu=\nu^{\prime}$. It is straightforward to see that any measurable $G$-equivariant factor $(Y^{\prime},\nu^{\prime})$ of a $(G,\mu)$-boundary $(Y,\nu)$ is still a $(G,\mu)$-boundary. In fact, the following Theorem states that there exists a pair $(B,\nu_{B})$ which is maximal and universal among $(G,\mu)$-boundaries.
Theorem 3.5 ([Fur73, Theorem 10.1]).
Given a locally compact second countable group $G$ with admissible probability measure $\mu$, there exists a maximal $(G,\mu)$-boundary, called the Poisson-Furstenberg boundary $(B,\nu_{B})$ of $(G,\mu)$, which is uniquely characterized by the following property:
Universality:
For every measurable $(G,\mu)$-boundary $(M,\nu)$, there is a $G$-equivariant measurable quotient map $p:(B,\nu_{B})\rightarrow(M,\nu)$, uniquely defined up to $\nu_{B}$-null sets.
The Poisson-Furstenberg boundary $B(G,\mu)$ associated to $\mu$ can also be defined as the space of ergodic components of the action of the shift map $S$ on $(\Omega,\operatorname{Haar}\otimes\mu^{\mathbb{N^{\ast}}})$. It is a measured space equipped by the push-forward measure $\nu_{B}$ of the measure $\mathbb{P}$ by the natural projection $\text{bnd}:\Omega\rightarrow B$. The space $(B,\nu_{B})$ is a $(G,\mu)$-space on which $G$ acts ergodically. In the next section, we state more of its ergodic properties.
3.2. Ergodic properties of strong boundaries
In this section, we review some notions of boundary theory that generalize the ergodic properties satisfied by the Poisson-Furstenberg boundary associated to a measured group $(G,\mu)$. We refer to [BF14] for the vocabulary of fiberwise isometric actions, and to [Duc13], [BDL16] for a more detailed introduction to this subject.
Throughout this section, $G$ is a locally compact second countable group. Recall that a $G$-space is a standard Lebesgue space $(S,\eta)$ endowed with a non-singular action of $G$, and a map $\psi:(S,\eta)\to(S^{\prime},\eta^{\prime})$ between $G$-spaces is called a $G$-map if it is a $G$-equivariant measurable map.
Definition 3.6.
Let $(S,\eta)$ be a $G$-space where $\eta\in\text{Prob}(S)$. We say that the $G$-action on $S$ is isometrically ergodic if for all separable metric space $(Y,d)$ on which $G$-acts by isometries, and for all $G$-equivariant measurable map $\psi:S\rightarrow Y$, alors $\psi$ is essentially constant. In other words, for any such $(Y,d)$, there is no non-trivial $G$-map
${f:(S,\eta)}$${(Y,d).}$$\scriptstyle{G}$
If the diagonal action of $G$ on $(S\times S,\eta\times\eta)$ is isometrically ergodic, we say that $G\curvearrowright S$ is doubly isometrically ergodic.
We can rephrase the previous definition in the following way. The $G$-action on $(S,\eta)$ is isometrically ergodic if for all metric space $(Y,d)$ on which $G$-acts by isometries,
$\text{Map}_{G}(S,Y)=\text{Map}_{G}(\{\ast\},Y)$
where we put the Borel $\sigma$-algebra on $(Y,d)$ given by the metric. Isometric ergodicity is a stronger property than ergodicity. Indeed, let $G\curvearrowright(S,\eta)$ be an isometrically ergodic action, and let $\psi:S\rightarrow[0,1]$ be a $G$-invariant measurable map. Take the trivial $G$-action on $[0,1]$ (so that $\psi$ is $G$-equivariant). By isometric ergodicity, $\psi$ is constant.
If $\mu$ is an admissible probability measure on $G$, then the $G$-action on its Poisson-Furstenberg boundary $(B(G,\mu),\nu)$ is doubly ergodic [Kai03, Theorem 3]. Actually, it satisfies even stronger ergodic properties, which we make precise now.
Let $\mathbb{X}$ and $A$ be $G$-spaces, and let $p:\mathbb{X}\rightarrow A$ be a $G$-map. We denote by $\mathbb{X}\times_{p}\mathbb{X}$ the fibered product over $p$, defined by $\mathbb{X}\times_{p}\mathbb{X}:=\{(x,y)\in\mathbb{X}\times\mathbb{X}\,|\,p(x)=p(y)\}$. Endow the fibered product $\mathbb{X}\times_{p}\mathbb{X}$ with the Borel structure from $\mathbb{X}\times\mathbb{X}$. The non-empty measurable sets $p^{-1}(a)$ for $a\in A$ are called the fibers of $p$, and we write $p^{-1}(a)=:\mathbb{X}_{a}\subseteq\mathbb{X}$.
Definition 3.7.
We say that a $G$-map $p:\mathbb{X}\rightarrow A$ admits a fiberwise isometric action of $G$ if there exists a Borel $G$-invariant map $d:\mathbb{X}\times_{p}\mathbb{X}\rightarrow\mathbb{R}$ such that for every fiber $\mathbb{X}_{a}\subseteq\mathbb{X}$, $d_{a}$ is a separable metric on $\mathbb{X}_{a}$. We call $d$ the fiberwise metric on $\mathbb{X}\times_{p}\mathbb{X}$.
By definition, the action of an element $g\in G$ on the fiber $\mathbb{X}\times_{p}\mathbb{X}$ induces isometries $g:\mathbb{X}_{a}\rightarrow\mathbb{X}_{ga}$: for $x,y\in\mathbb{X}_{a}$,
$d_{ga}(gx,gy)=d_{a}(x,y).$
Definition 3.8.
A measurable map $\pi:S\rightarrow Q$ between $G$-spaces is said to be relatively isometrically ergodic if for all fiberwise $G$-action on a $G$-map $p:\mathbb{X}\rightarrow A$ for which there exist $G$-equivariant maps $\psi:S\rightarrow\mathbb{X}$ and $\psi_{0}:Q\rightarrow A$ such that $p\circ\psi=\psi_{0}\circ\pi$, then there exists a relative section, i.e. a $G$-map $\phi:Q\rightarrow\mathbb{X}$ such that $\psi=\phi\circ\pi$ and $f_{0}=p\circ\phi$ almost everywhere. In other words, if the following diagram commutes
${S}$${\mathbb{X}}$${Q}$${A}$$\scriptstyle{\pi}$$\scriptstyle{\psi}$$\scriptstyle{p}$$\scriptstyle{\psi_{0}}$
then there exists an essentially well-defined $G$-map $\phi:Q\rightarrow\mathbb{X}$ such that the following diagram commutes.
${S}$${\mathbb{X}}$${Q}$${A}$$\scriptstyle{\pi}$$\scriptstyle{\psi}$$\scriptstyle{p}$$\scriptstyle{\psi_{0}}$$\scriptstyle{\phi}$
In the following, we might say that $\pi$ satisfies the relative isometric ergodicity (RIE) property.
Remark 3.9.
Let $(S,\eta)$ be a $G$-space such that the projection onto the first factor $\pi_{1}:S\times S\rightarrow S$ satisfies the RIE property. Then the $G$-action on $S$ is isometrically ergodic. Indeed, let $\psi:S\rightarrow(Y,d)$ be a $G$-map to a metric space $(Y,d)$ on which $G$ acts by isometries. Then $\tilde{\psi}:(b,b^{\prime})\in B\times B\mapsto\psi(b^{\prime})\in Y$ is a $G$-map and we can define the trivial fiberwise action
[…]
By RIE, there exists a $G$-map $\phi:S\rightarrow Y$ such that $(b,b^{\prime})$, $\phi(b)=f(b^{\prime})$ almost everywhere. By Fubini, it means that $\psi$ is essentially constant.
We say that $(A,\eta)$ is an amenable $G$-space if $(A,\eta)$ is a standard Borel space such that there exists a non-singular and amenable action $G\curvearrowright(A,\eta)$ in Zimmer’s sense, see [Zim84] and [Duc13, Section 7.1]. We do not give the original definition of an amenable action because actually, we will only use the following property.
Proposition 3.10.
Let $G$ be a locally compact second countable group, and let $(A,\eta)$ be an amenable $G$-space. Let $(X,d)$ be a compact metrizable space and $G\curvearrowright(X,d)$ an isometric action. Then there exists a measurable $G$-equivariant map $(A,\eta)\rightarrow\text{Prob}(X)$ which is essentially well-defined.
We are now ready for the definition of a $G$-boundary pair, introduced in [BF14].
Definition 3.11.
Let $(B_{-},\nu_{-})$ and $(B_{+},\nu_{+})$ be $G$-spaces. We say that $(B_{-},B_{+})$ is a $G$- boundary pair if:
  • the actions $G\curvearrowright B_{+}$ and $G\curvearrowright B_{-}$ are amenable;
  • the $G$-maps $B_{-}\times B_{+}\rightarrow B_{-}$ and $B_{-}\times B_{+}\rightarrow B_{+}$ are relatively metrically ergodic.
We say that a $G$-space $(B,\nu)$ is a $G$-boundary if $(B,B)$ is a $G$-boundary pair.
A reason why it is so convenient to use $G$-boundaries is that they give rigidity results on the group $G$, one of which is the existence of Furstenberg maps.
Theorem 3.12 ([BDL16, Theorem 1.1]).
Let $X$ be a complete CAT(0) space of finite telescopic dimension and let $G$ be a locally compact second countable group acting continuously by isometries on $X$ without invariant flats. If $(B,\nu)$ is a $G$-boundary, then there exists a measurable $G$-map $B\rightarrow\partial_{\infty}X$.
When $G$ is locally compact second countable, there always exists a $G$-boundary.
Theorem 3.13 ([BF14, Theorem 2.7]).
Let $G$ be a locally compact second countable group, and let $\mu$ be a spread-out and symmetric measure on $G$. Then the Poisson-Furstenberg boundary $(B(G,\mu),\nu_{B})$ is a $G$-boundary.
Corollary 3.14.
In particular, the $G$-action on $(B(G,\mu),\nu_{B})$ is metrically ergodic.
3.3. Metric fields and equivariant theorem
In the course of the proof, we will use the vocabulary of measurable metric fields, on which we can find accounts in [AH14] and in [Duc13]. This notion bears many similarities with the language of fiberwise isometric actions, as explained in [DLP23]. A good presentation of this notion is done in [BDL16].
3.3.1. Measurable metric fields
Let $(A,\eta)$ be a measurable space. A measurable metric field $\mathbf{X}$ over $A$ can be thought of as a way of attaching a metric space $X_{a}$ to each point $a\in A$ in a measurable way.
Definition 3.15.
Let $(A,\eta)$ be a Lebesgue space (a standard Borel space endowed with a Borel probability measure). Let $\mathbb{X}$ be a collection $\mathbb{X}=\{(X_{a},d_{a})\}_{a\in A}$ of complete separable metric spaces over $(A,\eta)$, and let $\{x^{n}\}_{n\in\mathbb{N}}$ be a countable set of elements of $\underset{a\in A}{\Pi}X_{a}$ such that:
  • $\forall n,m\in\mathbb{N},\,a\mapsto d_{a}(x^{n}_{a},x^{m}_{a})$ is measurable;
  • for $\eta$-almost every $a\in A$, $\{x^{n}_{a}\}_{n\in\mathbb{N}}$ is dense in $X_{A}$.
The family $\{x^{n}\}_{n\in\mathbb{N}}$ is called a fundamental family for $\mathbb{X}$. The data $((A,\eta),\mathbb{X},\{x^{n}\}_{n\in\mathbb{N}})$ is called a metric field over $(A,\eta)$.
A section $s:(A,\eta)\rightarrow\mathbb{X}$ is a map such that for every $a\in A$, $s(a)\in X_{a}$ and for every element $x^{n}$ of the fundamental family, $a\mapsto d_{a}(x^{n}_{a},s(a))$ is measurable.
Fix a locally compact second countable group $G$ endowed with its Haar measure, and let $(A,\eta)$ be a $G$-space. Now let $\mathbb{X}=\{X_{a}\}_{a\in A}$ be a metric field over $A$. We say that $G$ acts on $\mathbb{X}$ by the cocycle $\alpha$ if:
  • for all $g\in G$ and almost every $a\in A$, $\alpha(g,a)\in\operatorname{Isom}(X_{a},X_{ga})$;
  • for all $g,g^{\prime}\in G$ and almost every $a\in A$, $\alpha(gg^{\prime},a)=\alpha(g,g^{\prime}a)\alpha(g^{\prime},a)$;
  • for all element $x^{n},x^{m}$ of the fundamental family associated to $\mathbb{X}$, the map $(g,a)\mapsto d_{a}(x^{n}_{a},\alpha(g,g^{-1}a)x^{m}_{g^{-1}a})$ is measurable.
A section $s:(A,\eta)\rightarrow\{X_{a}\}_{a\in A}$ is then said to be $\alpha$-invariant if for all $g\in G$ and almost every $a\in A$, $s(ga)=\alpha(g,a)s(a)$.
Definition 3.16.
Let $\mathbb{X}=\{X_{a}\}_{a\in A}$ be a metric field over $(A,\eta)$. We say that $\mathbb{X}$ is a CAT($\kappa$)-field if for almost every $a$, $X_{a}$ is a $\text{CAT}(\kappa)$ space. Similarly, if for almost every $a$, $X_{a}$ is a Euclidean space, we say that $\mathbb{X}$ is a Euclidean field. A subfield of a CAT(0) field is a collection $\{Y_{a}\}_{a}$ of non-empty closed convex subsets such that for every section $x$ of $\mathbb{X}$, the function $a\mapsto d(x_{a},Y_{a})$ is measurable. A subfield is called invariant if for all $g$, and $\eta$-almost every $a\in A$,
$\alpha(g,g^{-1}a)Y_{g^{-1}a}=Y_{a}.$
Let $\mathbb{X}$ be a CAT(0)-field over $(A,\eta)$. There is a way to define the boundary field $\partial\mathbb{X}$ of $\mathbb{X}$, but one must be careful with the measurable structure and the separability conditions. For instance, the boundary of the hyperbolic plane $\mathbb{H}^{2}$, endowed with the Tits metric, is an uncountable discrete space. However, it is possible if one considers the compactification by horofunctions, as explained in [Duc13, Section 9.2].
Let $G$ be a locally compact second countable group and $B$ an ergodic $G$-space. Let $\mathbb{X}$ be a measurable field of CAT(0) spaces of finite telescopic dimension endowed with a $G$-cocycle $\alpha:G\times B\rightarrow\operatorname{Isom}(\mathbb{X})$. Fix a fundamental family $\{x^{n}\}$ for $\mathbb{X}$. For $b\in B$, let $C_{b}$ be the locally convex linear space of convex functions that vanish at $x^{0}_{b}$, and endow $C_{b}$ with the topology of pointwise convergence. Let $L_{b}$ be the subset of $C_{b}$ which consists of functions that are moreover 1-Lipschitz. A metric on $L_{b}$ is given by
$D_{b}(f,g)=\sum\frac{|f(x^{i}_{b})-g(x^{i}_{b})|}{2^{i}d_{b}(x^{i}_{b},x^{0}_{b})}.$
Let $\iota_{b}:X_{b}\to L_{b}$ be the map defined by
$\iota_{b}(x):y\in X_{b}\mapsto d_{b}(x,y)-d(x,x^{0}_{b}),$
which is indeed a 1-Lipschitz map on $X_{b}$. Define $K_{b}$ to be the closed convex hull of $\iota(X_{b})$ in $L_{b}$. Then $\mathbb{K}:=(K_{b})_{b}$ is a measurable field of compact spaces over $B$, with a fundamental family given by $\{\iota_{b}(x^{i}_{b})\}_{i}$. The metric field $\mathbb{K}$ corresponds to the horocompactification of the metric field $\mathbb{X}$.
Let $\mathbb{X}$ is a measurable field of proper metric spaces endowed with a $G$-cocycle $\alpha:G\times B\rightarrow\operatorname{Isom}(\mathbb{X})$. Fix a fundamental family $\{x^{n}\}$ for $\mathbb{X}$. Then there exists a measurable structure on the collection $\text{Prob}(\mathbb{X}):=\{\text{Prob}(X_{b})\}_{b}$, coming from a Borel structure on the set $C(\mathbb{X}):=\{C(X_{b})\}_{b}$ of Banach spaces of continuous functions over $\mathbb{X}$. We refer to [AH14, Theorem 2.19] for details.
The link between measurable metric fields and fiberwise isometric actions is given by the following Lemma.
Lemma 3.17 ([DLP23, Lemma 4.11]).
Let $G$ be a countable discrete group and $\mathbb{X}=\{(X_{a},d_{a})\}_{a\in A}$ be a measurable metric field over a Lebesgue $G$-space $(A,\eta)$, and let us assume that $G$ acts on $\mathbb{X}$ with some cocycle $\alpha$. Then there is a $G$-invariant Borel subset $A_{0}\subseteq A$ of full $\eta$-measure, a standard Borel structure on $\mathbb{X}:=\bigsqcup_{a\in A_{0}}X_{a}$, and a Borel map $p:\mathbb{X}\rightarrow A_{0}$ such that $p$ admits a $G$-fiberwise isometric action. Moreover, the fiber $p^{-1}(a)$ is $X_{a}$ with the metric $d_{a}$.
If $s$ is an invariant section of $\mathbb{X}$, then $s$ corresponds canonically to a $G$-equivariant measurable map $A_{0}\rightarrow\mathbb{X}$.
3.3.2. Adams-Ballmann equivariant theorem for amenable actions
Recall that a flat is a closed convex subspace which is isometric to some finite dimensional Euclidean space, and that a point is a flat of dimension 00. In [AB98, Theorem 1], Adams and Ballmann prove a general dichotomy for amenable groups acting on Hadamard spaces.
Theorem 3.18 (Adams-Ballmann Theorem).
Let $X$ be a locally compact complete CAT(0) space. If $G$ is an amenable group of isometries of $X$, then at least one of the following proposition hold:
  1. (i)
    $G$ fixes a point in $\partial_{\infty}X$;
  2. (ii)
    $X$ contains a $G$-invariant flat.
The following result is a (measurable) version of the equivariant Adams-Ballmann Theorem, but for amenable actions, and was proven by Duchesne.
Theorem 3.19 ([Duc13, Theorem 1.8]).
Let $G$ be a locally compact second countable group and $B$ an amenable ergodic $G$-space. Let $\mathbb{X}$ be a measurable metric field of Hadamard spaces of finite telescopic dimension over $B$. If $G$ acts on $\mathbb{X}$, then either there is an invariant section of the boundary field $\partial\mathbb{X}$, or there is an invariant Euclidean subfield of $\mathbb{X}$.
In Section 5, we will use this correspondence to study the properties of $G$-equivariant measurable maps $B\rightarrow X^{\infty}$.
In the course of the proof, we will use the fact that taking the circumcenters in a subfield of bounded closed convex subsets gives a measurable map, see [Duc13, Lemma 8.7] for the general complete case.
Proposition 3.20 ([AH14, Lemma 3.2]).
Let $\mathbb{X}$ be a metric field of proper CAT(0) spaces over a Borel space $(A,\eta)$, and let $\{B_{a}\}_{a\in A}$ be a Borel subfield of bounded subsets of $\mathbb{X}$. Then the circumradius function $a\in A\mapsto r_{a}(B_{a})$ is Borel, and the section of circumcenters
$\text{circ}:\mathbb{B}$$\rightarrow$ $\mathbb{X}$
$B_{a}$ $\mapsto$ $\text{circ}(B_{a}),$
where $\text{circ}(B_{a})$ is the circumcenter of $B_{a}$, is Borel.
4. Measurable structures and barycenters
Let $X$ be a complete separable thick building of type $\tilde{A}_{2}$ and let $G$ be a locally compact second countable group acting continuously by isometries on $X$ without fixed points nor invariant flats. Up to taking a subgroup of index 2, we may assume that $G$ preserves the types of the vertices at infinity. In this Section, we give an explicit construction of some measurable structures and useful equivariant maps so that the results of Section 3.2 can apply.
4.1. Measurable structures on the panel trees
Recall that for every $u\in X^{\infty}_{1}$, $T_{u}$ is the panel tree associated to the vertex $u$. Endow $T_{u}$ with the separable metric $d_{u}$ given by Equation 3.
Lemma 4.1.
The collection of panel trees $T_{1}=\bigsqcup_{u}T_{u}$ admits a standard Borel structure and the natural projection $p:T_{1}\rightarrow X^{\infty}_{1}$ admits a $G$-fiberwise isometric action.
Proof.
The collection $\{(T_{u},d_{u})\}_{u\in X^{\infty}_{1}}$ is a family of complete, separable metric spaces. Let us construct a fundamental family for $\{(T_{u},d_{u})\}_{u}$ over $X^{\infty}_{1}$. Let $\{y_{n}\}_{n}$ be a countable dense set of special vertices of the building $X$. For every $n$, and every $u\in X^{\infty}_{1}$, consider
$x^{n}_{u}:=\pi_{u}(y_{n})$
where $\pi_{u}(y_{n})$ is the class (for the strong asymptote relation) of the geodesic ray based at $y$ in the direction of $u$. Observe that for every $u$, $\{x^{n}_{u}\}_{n}$ is dense in $T_{u}$. Fix $n,m\in\mathbb{N}$. Then the function
$u\in X^{\infty}_{1}\mapsto d_{u}(x^{n}_{u},x^{m}_{u})$
is measurable. Indeed, for every $y\in X$, the application $\xi\in\partial_{\infty}X\mapsto\gamma_{y}^{\xi}$ is continuous for the topology of uniform convergence on compact sets, where $\gamma_{y}^{\xi}$ is the unique geodesic ray based at $y$ representing $\xi$. Recall that by definition,
[…]
Since we consider the infimum of measurable functions, we have that for all $n,m$, the map $u\in X^{\infty}_{1}\mapsto d_{u}(x^{n}_{u},x^{m}_{u})$ is then measurable. Therefore, the collection $\{x^{n}\}$ is a fundamental family for the metric field $\{T_{u}\}_{u}$ over $X^{\infty}_{1}$.
Let us now prove that $G$ acts on the metric field $\{T_{u}\}_{u}$. If $u\in X^{\infty}_{1}$ is a vertex at infinity and $\xi\in T_{u}$ is a class (for the strong asymptote relation) of geodesic rays parallel to $u$, the element $g\in G$ acts as a cocycle $\alpha$ with
$\alpha(g,u)(u,\xi)=g\cdot(u,\xi)=(gu,g\xi),$
where $g\xi$ is the class (for the strong asymptote relation) of $g\gamma$, for any geodesic ray $\gamma$ in the class of $\xi$. It is clear for all $g\in G$,
$\alpha(g,u)\in\operatorname{Isom}(T_{u},T_{gu}).$
It is clear that $\alpha$ satisfies the cocycle relation: for all $g,g^{\prime}\in G$, $u\in X^{\infty}_{1}$, $\xi\in T_{u}$,
$g^{\prime}\cdot(g\cdot(u,\xi))=(g^{\prime}g)\cdot(u,\xi).$
It remains to prove that $G$ acts on this metric field in a measurable way.
Take $n,m\in\mathbb{N}$. Observe that for $g\in G$, $u\in X^{\infty}_{1}$,
$g\cdot x^{m}_{g^{-1}u}$ $=$ $g\cdot\pi_{g^{-1}u}(y^{m})$
$=$ $g\cdot[Q(y^{m},g^{-1}u)]$
$=$ $[Q(gy^{m},u)]\in T_{u}.$
As $G$ acts continuously on $X$ and preserves the type of the vertices at infinity, we obtain that
$(g,u)\mapsto d_{u}(x^{n}_{u},g\cdot x^{m}_{g^{-1}u})$
is measurable.
Now by Lemma 3.17, there is a standard Borel structure on $T_{1}:=\bigsqcup_{u\in X^{\infty}_{1}}T_{u}$, and a Borel map $p:T_{1}\rightarrow X^{\infty}_{1}$ such that $p$ admits a $G$-fiberwise isometric action and such that $p^{-1}(u)=T_{u}$ for every $u\in X^{\infty}_{1}$. ∎
Recall that we denote by $\partial T_{1}$ the union of all the boundaries of the panel trees $T_{u},u\in X^{\infty}_{1}$. Namely,
$\partial T_{1}=\{(u,C)\,|\,u\in X^{\infty}_{1},C\in\partial T_{u}\}.$
For $(u,\xi)\in T_{1}$, that is, $u\in X^{\infty}_{1}$ and $\xi\in T_{u}$, we can define the following metric on $\partial T_{u}$:
$d_{u,\xi}(C,D):=\exp(-(C|D)_{\xi}),$
with $(C,D)\in\partial T_{u}$, and where (.|.$(.\,|\,.)_{\xi}$ is the Gromov product on the tree $T_{u}$ based at $\xi\in T_{u}$.
The next Lemma states that the natural projection $T_{1}\times\partial T_{1}\rightarrow T_{1}$ admits a $G$-fiberwise isometric action for this metric.
Lemma 4.2.
The collection $T_{1}\times\partial T_{1}=\bigsqcup_{u\in X^{\infty}_{1}}T_{u}\times\partial T_{u}$ admits a standard Borel structure and the natural projection $q:T_{1}\times\partial T_{1}\rightarrow T_{1}$ admits a $G$-fiberwise isometric action.
Proof.
Endow $T_{1}$ with the measurable structure given by Lemma 4.1. Consider the family $\{T_{u}\times\partial T_{u}\}_{u\in X^{\infty}_{1}}$, and call $q$ the projection on the first coordinate
$q$ ::: $\{T_{u}\times\partial T_{u}\}_{u\in X^{\infty}_{1}}\longrightarrow T_{1}$
$(u,\xi,C)\longmapsto(u,\xi)\in T_{u}.$
For every $u\in X^{\infty}_{1}$, $\xi\in T_{u}$, the metric $d_{u,\xi}$ on the fiber $q^{-1}\{(u,\xi)\}=\{(u,\xi)\}\times\partial T_{u}$ is complete and separable. For convenience, we denote the metric spaces $(q^{-1}\{(u,\xi)\},d_{u,\xi})$ by $\partial T_{u,\xi}$. In order to prove that this is a metric field over $T_{1}$, we need to construct a fundamental family. Consider a countable dense family $\{C_{n}\}$ of chambers in $\text{ch}(X^{\infty})$. For every $u\in X^{\infty}_{1}$, define
$C^{n}_{u}:=\text{proj}_{u}(C_{n})\in\text{Res}(u).$
The projection onto the vertex $u$ is continuous, see [CMR20, Proposition 3.14]. Moreover, if one fixes $C\in\text{ch}(X^{\infty})$, the application $u\in X^{\infty}_{1}\mapsto\text{proj}_{u}(C)$ is continuous. By Proposition 2.7, $C^{n}_{u}\in\partial T_{u}$. Consequently, for every $(u,\xi)\in T_{1}$, define
$\mathcal{C}^{n}_{u,\xi}:=(u,\xi,C^{n}_{u})\in T_{u}\times\partial T_{u}.$
For all $(u,\xi)$, the collection $\{\mathcal{C}^{n}_{u,\xi}\}_{n}$ is dense in the fiber $\partial T_{u,\xi}$. Fix $n,m\in\mathbb{N}$, and let $(u,\xi)\in T_{1}$. We have that
$d_{u,\xi}(\mathcal{C}^{n}_{u,\xi},\mathcal{C}^{m}_{u,\xi})$ $=$ $d_{u,\xi}(C^{n}_{u},C^{m}_{u})$
$=$ $\exp(-(C^{n}_{u}|C^{m}_{u})_{\xi})$
$=$ $\exp(-(\text{proj}_{u}(C_{n})|\text{proj}_{u}(C^{m}))_{\xi}).$
By composition, the application
$(u,\xi)\in T_{1}\longmapsto d_{u,\xi}(\mathcal{C}^{n}_{u,\xi},\mathcal{C}^{m}_{u,\xi})$
is measurable. The collection $\{\mathcal{C}^{n}_{u,\xi}\}_{n}$ is then a fundamental family for the metric field $\{T_{u}\times\partial T_{u}\}_{u\in X^{\infty}_{1}}$ over $T_{1}$.
By Lemma 4.1, the projection $T_{1}\rightarrow X^{\infty}_{1}$ admits a $G$-fiberwise isometric action. For $g\in G$, and $(u,\xi,C)\in\{T_{u}\times\partial T_{u}\}$,
$\alpha(g,(u,\xi))(u,\xi,C)=g\cdot(u,\xi,C)=(gu,g\xi,gC),$
where $(gu,g\xi)$ was defined in the proof of Lemma 4.1, and $gC\in\partial T_{gu}$ is the boundary point of $\partial T_{gu}$ that corresponds to the chamber $gC\in\text{Res}(gu)$ by Proposition 2.7. With this action, for all $g\in G$,
$\alpha(g,(u,\xi))\in\operatorname{Isom}(\partial T_{u,\xi},\partial T_{g\cdot(u,\xi)}).$
Moreover, the projection $q$ is $G$-equivariant. Last, for every $n,m\in\mathbb{N}$, the measurability of
$(u,\xi)\in T_{1}\longmapsto d_{u,\xi}(\mathcal{C}^{n}_{u,\xi},g\cdot\mathcal{C}^{m}_{g^{-1}\cdot(u,\xi)})$
comes from the continuity of the $G$-action on the chambers $\text{ch}(X^{\infty})$. This proves that $G$ acts on the metric field $\{T_{u}\times\partial T_{u}\}_{u\in X^{\infty}_{1}}$ with a natural cocycle. By Lemma 3.17, there is a standard Borel structure on $T_{1}\times\partial T_{1}=\bigsqcup_{u}T_{u}\times\partial T_{u}$ such that the projection $T_{1}\times\partial T_{1}\rightarrow T_{1}$ is Borel and admits a $G$-fiberwise isometric action. ∎
4.2. Barycenter map
Consider the set $(\partial T_{1})^{2}$ defined by
$(\partial T_{1})^{2}=\{(u,C_{1},C_{2}),u\in X^{\infty}_{1},C_{1},C_{2}\in\partial T_{u}\}.$
Denote by $\mathcal{D}(\partial T_{1})$ the diagonal:
$\mathcal{D}(\partial T_{1})=\{(u,C_{1},C_{2}),u\in X^{\infty}_{1},C_{1}=C_{2}\in\partial T_{u}\}\subseteq(\partial T_{1})^{2}.$
Define the application:
$\tau$ ::: $(\partial T_{1})^{2}\to(\partial T_{1})^{2}$
$(u,C_{1},C_{2})\in(\partial T_{u})^{2}\mapsto(u,C_{2},C_{1})\in(\partial T_{u})^{2}.$
Let $(\partial T_{1})^{(2)}$ be the measurable set $(\partial T_{1})^{2}-\mathcal{D}(\partial T_{1})$, and define the space $\partial_{2}T_{1}$ as the set $(\partial T_{1})^{(2)}/\tau$ of unordered pairs of distinct points of $\partial T_{1}$. Similarly, we define by $\partial_{n}T_{1}$ as the set of unordered $n$-tuples of points of $\partial T_{1}$ that are pairwise distinct. Endow $\partial_{2}T_{1}$ and $\partial_{n}T_{1}$ with the product measurable structure coming from $\partial T_{1}$.
In the course of the proof, we will need the existence of a well-defined barycenter map on the set $\partial_{3}T_{1}$.
Proposition 4.3.
For all $n\geq 3$, there exists a measurable $G$-equivariant map $\partial_{n}T_{1}\to T_{1}$.
Proof.
Let $u\in X^{\infty}_{1}$. Let $S\subseteq\partial_{n}T_{u}$ be a set of $n$ distinct points in $\partial T_{u}$. Consider the function
$F_{S}:x\in T_{u}\mapsto\sum_{C\neq C^{\prime}\in S}(C|C^{\prime})_{x}.$
It is standard to check that the function $F_{S}$ is proper and convex, see for example [BM96, Section 2]. Therefore its minimum $\min(F_{S})=\{x\in T_{u}\;|\;F_{S}(x)=\inf F_{S}\}$ is non-empty, bounded and closed. Since $T_{u}$ is a tree, we can take the circumcenter of this set, which gives a continuous map
$\text{bar}_{u}:\partial_{n}T_{u}\to T_{u}.$
We define
bar ::: $\partial_{n}T_{1}\to T_{1}$
$(u,C_{1},\dots,C_{n})\in\partial_{n}T_{u}\mapsto\text{bar}_{u}(u,C_{1},\dots,C_{n})\in T_{u}.$
Let $G$ act on $\partial_{n}T_{1}$ with the natural diagonal action. Then the map bar is $G$-equivariant:
$\text{bar}(g\cdot(u,C_{1},\dots,C_{n}))$ $=$ $\text{bar}_{gu}(g\cdot(u,C_{1},\dots,C_{n}))$
$=$ $g\cdot\text{bar}(u,C_{1},\dots,C_{n}).$
Moreover, bar is measurable for the measurable structures defined by Lemma 4.1 and 4.2. Indeed, if we take a measurable section $s:X^{\infty}_{1}\to\partial_{n}T_{1}$, then the map
$u\in X^{\infty}_{1}\mapsto\text{bar}_{u}(s(u))$
is the composition of two maps. First, the collection $\{\min(F_{s(u)})\}_{u\in X^{\infty}_{1}}$ is a Borel subfield of non-empty closed convex bounded subsets of the metric field $T_{1}$, in the sense of Definition 3.16. Indeed, for any other section $s^{\prime}$,
$u\in X^{\infty}_{1}\mapsto d_{u}(s^{\prime}(u),\min(F_{s(u)}))$
is measurable by the definition of $u\mapsto F_{s(u)}$. Next, by Proposition 3.20, taking the section of circumcenters gives a Borel map
$\{\min(F_{s(u)})\}_{u\in X^{\infty}_{1}}\rightarrow T_{u}.$
This ends the proof. ∎
Let $Y$ be a proper $\text{CAT}(-1)$ space, and let $\text{Prob}_{3}(\partial_{\text{Grom}}Y)$ be the space of positive measures on $\partial_{\text{Grom}}Y$ whose support contains at least three points. In [BM96, Proposition 2.1], Burger and Mozes prove that there exists a $\operatorname{Isom}(Y)$-equivariant map
(6) $\text{bar}:\text{Prob}_{3}(\partial_{\text{Grom}}Y)\to Y$
whose restriction to any $\operatorname{Isom}(Y)$-orbit is continuous. We shall now prove that there exists a measurable version of this fact, for the metric field $T_{1}$.
Let $u\in X^{\infty}_{1}$ be a vertex at infinity of type 1 and let $\text{Prob}_{3}(\partial T_{u})$ be the space of positive measures on $\partial T_{u}$ whose support contains at least three points.
Lemma 4.4.
There exists a measurable map $\varphi_{u}:\text{Prob}_{3}(\partial T_{u})\rightarrow\text{Prob}(T_{u})$.
Proof.
Let $\nu\in\text{Prob}_{3}(\partial T_{u})$. For the moment, assume that $\nu$ is non-atomic. Then consider the probability measure $\tilde{\nu}:=\nu\otimes\nu\otimes\nu$ on $(\partial T_{u})^{3}$. Let $\mathcal{D}^{3}(\partial T_{u})$ be the space of triples $(u,\xi_{1},\xi_{2},\xi_{3})\in(\partial T_{u})^{3}$ such that at least two entries coincide. The set $\mathcal{D}^{3}(\partial T_{u})$ is measurable. Since $\nu$ is non-atomic, $\tilde{\nu}$ gives zero measure to $\mathcal{D}^{3}(\partial T_{u})$ and we can see $\tilde{\nu}$ as a measure on $(\partial T_{1})^{(3)}$. Passing to the quotient, we can see $\tilde{\nu}$ as a measure on $\partial_{3}T_{1}$, which we do without changing the notation. Now, thanks to the barycenter map defined in Proposition 4.3, $\varphi_{u}(\nu):=(\text{bar}_{u})_{\ast}\tilde{\nu}$ is a probability measure on $T_{u}$.
Assume now that $\nu$ is purely atomic, that is, for every $\xi\in\operatorname{\text{supp}}(\nu)$, $\nu(\xi)>0$. Since the support of $\nu$ is of cardinal at least $3$, $\nu$ has at least 3 atoms $\xi_{1},\xi_{2},\xi_{3}$. Again, let $\tilde{\nu}:=\nu\otimes\nu\otimes\nu$ be the product probability measure on $(\partial T_{u})^{3}$. By definition, $\tilde{\nu}(\xi_{1},\xi_{2},\xi_{3})>0$ and in particular $\tilde{\nu}((\partial T_{1})^{(3)})>0$. Recall that taking the restriction of a measure to an open subset is a measurable operation. We can then consider
$\bar{\nu}=\frac{1}{\tilde{\nu}((\partial T_{1})^{(3)})}\tilde{\nu}_{|(\partial T_{1})^{(3)}},$
which is a probability measure on the set of triples of pairwise distinct points in $\partial T_{u}$. Again, we can see $\bar{\nu}$ as a probability measure on $\partial_{3}T_{1}$, and applying the barycenter map $\text{bar}_{u}$ from Proposition 4.3, one obtains a probability measure $\varphi_{u}(\nu)$ on $T_{u}$.
If now $\nu\in\text{Prob}_{3}(\partial T_{u})$ is any probability measure, we can decompose $\nu$ as a sum $\nu=\nu_{c}+\nu_{a}$, where $\nu_{c}$ is non-atomic and $\nu_{a}$ is purely atomic. By the previous case, we can assume that the non-atomic part is nonzero $\nu_{c}>0$. Therefore, the probability measure $\nu_{c}/\nu_{c}(\partial T_{u})$ is non-atomic, and by the first case we can measurably associate a probability measure on $T_{u}$. ∎
We denote by $\text{Prob}_{3}(\partial T_{1})$ the subfield of $\text{Prob}(\partial T_{1})$ of measurable sections $u\in X^{\infty}_{1}\mapsto\nu_{u}\in\text{Prob}_{3}(\partial T_{u})$ for the measurable structure on $\partial T_{1}$.
Lemma 4.5.
There exists a measurable and $G$-equivariant map $\varphi:\text{Prob}_{3}(\partial T_{1})\rightarrow\text{Prob}(T_{1})$.
Proof.
On every fiber $u\in X^{\infty}_{1}$, the application $\varphi_{u}:\text{Prob}_{3}(\partial T_{u})\rightarrow\text{Prob}(T_{u})$ is measurable. Moreover, every operation that we did in the proof of Lemma 4.4 was $G$-equivariant, i.e. if $\nu=\{\nu_{u}\}_{u\in X^{\infty}_{1}}\in\text{Prob}_{3}(\partial T_{1})$, then for every $u\in X^{\infty}_{1}$,
$g\cdot(\varphi_{u}(\nu_{u}))=\varphi_{gu}(\nu_{gu}).$
Measurability of this map comes from Proposition 4.3 and the operations in the proof of Lemma 4.4. This proves the measurability and $G$-equivariance of the application. ∎
For self-containment, we now recall the classical construction of a barycenter map for the set of measures on a tree. Let $\varepsilon<\frac{1}{2}$, $\nu\in\text{Prob}(T_{u})$ and $\xi\in T_{u}$. For any $R>0$, define
$F^{\nu}_{\xi}(R):=\nu(B(\xi,R)),$
where $B(\xi,R)$ is the ball of radius $R$ and center $\xi$ in $T_{u}$ for the metric $d_{u}$ given by equation (3). It is clear that $F^{\nu}_{\xi}(R)\to 1$ as $R\to\infty$, hence for all $\xi\in T_{u}$, there exists a well-defined and finite
$R^{\nu}_{\xi,\varepsilon}:=\inf\{R\;|\;F^{\nu}_{\xi}(R)>1-\varepsilon\}.$
Also, denote by $R^{\nu}_{\varepsilon}:=\inf_{\xi\in T_{u}}R^{\nu}_{\xi,\varepsilon}$. We can then define the measurable application
$\tilde{\beta}^{u}_{\varepsilon}(\nu):=\overline{\{\xi\in T_{u}\;|\;R^{\nu}_{\xi,\varepsilon}<R^{\nu}_{\varepsilon}+1\}}.$
Check that $\tilde{\beta}^{u}_{\varepsilon}(\nu)$ is a closed bounded set on $T_{u}$. Indeed, if $\xi_{1},\xi_{2}\in\tilde{\beta}^{u}_{\varepsilon}(\nu)$, then by definition
$\nu(\xi\in T_{u}\;|\;B(\xi_{i},R^{\nu}_{\varepsilon}+1))>1/2\text{ for $i=1,2$},$
and then $B(\xi_{1},R^{\nu}_{\varepsilon}+1))\cap B(\xi_{2},R^{\nu}_{\varepsilon}+1))\neq\emptyset$. Now since trees are CAT(0) spaces, Proposition 3.20 states that we can measurably associate to each bounded set its circumcenter. We then have a measurable application
$\beta^{u}_{\varepsilon}:\nu\in\text{Prob}(T_{u})\mapsto\text{circ}(\tilde{\beta}^{u}_{\varepsilon}(\nu))\in T_{u}.$
From the definition, $\beta_{\varepsilon}^{u}$ is $\operatorname{Isom}(T_{u})$-equivariant.
Considering the metric field $\text{Prob}(T_{1})$ over $X^{\infty}_{1}$, what we have obtained is that the map
$\beta_{\varepsilon}:\text{Prob}(T_{1})$$\to$ $T_{1}$
$\{\nu_{u}\}_{u\in X^{\infty}_{1}}$ $\mapsto$ $\{\beta^{u}_{\varepsilon}(\nu_{u})\}_{u\in X^{\infty}_{1}}$
is well-defined, measurable, and $G$-equivariant for the natural $G$-action on $\text{Prob}(T_{1})$. Composing $\beta_{\varepsilon}$ and $\varphi$ from Lemma 4.5, we have obtained the following result.
Proposition 4.6.
There exists a measurable and $G$-equivariant map $\text{Prob}_{3}(\partial T_{1})\rightarrow T_{1}$.
We define $T_{2}$ as the collection of panel trees over the vertices at infinity $X^{\infty}_{2}$ of type 2, and we proceed similarly for $\partial T_{2}$ and $\partial_{i}T_{2}$. All the results in this section apply for these metric fields in the same way.
5. Boundary maps in $\tilde{A}_{2}$-buildings
Let $X$ be a thick building of type $\tilde{A}_{2}$ which is metrically complete and separable, and let $G$ be a locally compact second countable group acting continuously by automorphisms on $X$ without fixed points nor invariant flats. For the rest of this chapter, let $(B,\nu)$ be a $G$-boundary. By Theorem 3.12, there exists a measurable $G$-map $\psi:B\rightarrow\partial_{\infty}X$.
Recall that up to taking a subgroup of index 2, we assume that the $G$-action is type preserving, so that $G$ acts on $X^{\infty}_{1}$ and on $X^{\infty}_{2}$. Moreover, the partition of the visual boundary $\partial_{\infty}X$ between regular points (that is, classes of geodesic rays whose endpoints belong to the interiors of the chambers at infinity), and singular points (that is, classes of geodesic rays that represent the vertices at infinity) is measurable because $X^{\infty}_{1}$ and $X^{\infty}_{2}$ are closed in $\partial_{\infty}X$. In other words, we have the measurable decomposition $\partial_{\infty}X=X^{\infty}_{1}\sqcup X^{\infty}_{2}\sqcup(\partial_{\infty}X-X^{\infty}_{1}\cup X^{\infty}_{2})$, and the set of regular points $\partial_{\infty}X-X^{\infty}_{1}\cup X^{\infty}_{2}$ projects equivariantly and measurably onto $\text{ch}(X^{\infty})$.
In order to study stationary measures on $\text{ch}(X^{\infty})$, one needs to characterize the $G$-maps $B\rightarrow\text{Prob}(\text{ch}(X^{\infty}))$. Actually, we will prove that there exists a unique boundary map $\psi B\to\text{ch}(X^{\infty})$, and next that every stationary measure disintegrates according to $\psi$. The purpose of this section is then to study boundary maps. First, we prove that such a map exists.
Theorem 5.1.
Let $X$ be a thick, metrically complete and separable building of type $\tilde{A}_{2}$, $G$ be a locally compact second countable group and $G\curvearrowright X$ a non-elementary action by type-preserving isometries. Let $(B,\nu)$ be a $G$-boundary. Then there exists a measurable map $B\rightarrow\text{ch}(X^{\infty})$ which is $G$-equivariant.
The proof of Theorem 5.1 is made in Section 5.1 below.
5.1. A trichotomy for the boundary map
Thanks to Theorem 3.12, there exists a Furstenberg map $\psi:B\to\partial_{\infty}X$. The purpose of this section is to investigate the characteristics of this map, using the nice ergodic properties of $B$ and the geometric structure of $\partial_{\infty}X$. If almost surely, $\psi(b)$ is in a chamber of the building at infinity, we get a $G$-map $B\to\text{ch}(X^{\infty})$, which is what we want. This leads us to study the case where the target space is the set of vertices at infinity. In the following, we might refer to $G$-equivariant measurable maps simply as $G$-maps.
Lemma 5.2.
Consider the same assumptions as in Theorem 5.1, and assume that there exists a $G$-map $B\rightarrow X^{\infty}_{1}$. Then there is either a $G$-map $B\rightarrow T_{1}$, or a $G$-map $B\rightarrow\partial T_{1}$, or a $G$-map $B\rightarrow\partial_{2}T_{1}$.
Proof.
Let $\phi$ be a $G$-map $\phi:B\rightarrow X^{\infty}_{1}$, then by Lemma 4.1, the panel trees $\{T_{\phi(b)}\}_{b\in B}$ endowed with their natural metrics form a measurable metric field over $B$. Moreover, there is a natural $G$-action on $T_{1}$ over $X^{\infty}_{1}$. Each individual $T_{u}$ is CAT(0), complete and of finite telescopic dimension, so we can apply the measurable equivariant Theorem 3.19 to the metric field $T_{1}$ over $X^{\infty}_{1}$. Therefore either there exists an invariant section of the boundary field $\partial T_{1}$, or there is an invariant Euclidean subfield of $T_{1}$. Note that by Lemma 3.17, an invariant section of the boundary field gives a measurable $G$-map
$B\rightarrow\partial T_{1}.$
Otherwise, there is an invariant Euclidean subfield. If this subfield is made of closed bounded subfields in $T_{1}$, then taking the section of circumcenters as in Proposition 3.20 yields a $G$-map $B\rightarrow T_{1}$. If this subfield consists of lines, it gives a $G$-map $B\rightarrow\partial_{2}T_{1}$. ∎
Now our goal is to rule out two of the three previous possibilities.
Lemma 5.3.
Under the same assumptions as in Theorem 5.1, there is no $G$-map $B\rightarrow T_{1}$.
Proof.
Let us assume that there is a $G$-map $\psi:B\rightarrow T_{1}$. Consider the natural projection $p:T_{1}\rightarrow X^{\infty}_{1}$, and denote $\psi_{1}=p\circ\psi:B\rightarrow X^{\infty}_{1}$. By Lemma 4.1, $p$ is measurable and $G$-equivariant, hence by composition so is $\psi_{1}$. Take $u,v\in X^{\infty}_{1}$ two vertices at infinity of type $1$. Recall that for the boundary of an $\tilde{A}_{2}$-building, if $u\neq v$ are both of the same type, the projection $\text{proj}_{u}(v)$ is a chamber in the residue $\text{Res}(u)$. This map is continuous in $v$, and measurable in $u$. Thanks to Proposition 2.7, we can identify the chamber $\text{proj}_{u}(v)$ to an endpoint of $\partial T_{u}$, which we do without changing the notation.
Observe that because $\psi_{1}$ is $G$-equivariant, the set $\{(b,b^{\prime})\;|\;\psi_{1}(b)=\psi_{1}(b^{\prime})\}$ is $G$-invariant. Since the action of $G$ on $B\times B$ is doubly ergodic, this set is either null or conull. If $\nu\otimes\nu$ almost surely, $\psi_{1}(b)=\psi_{1}(b^{\prime})$, then varying $b$ and $b^{\prime}$ separately and applying Fubini gives that $\psi_{1}$ is almost surely constant. But since $\psi_{1}$ is $G$-equivariant, it means that there is a $G$-fixed point in $\partial_{\infty}X$, which is forbidden by non-elementarity of the $G$-action. As a consequence, $\psi_{1}(b)\neq\psi_{1}(b^{\prime})$ almost surely.
This discussion allows us to define the map:
ϕ::italic-$\phi:$ $B\times B$ $\longrightarrow\partial T_{1}$
$(b,b^{\prime})$ $\longmapsto(\psi_{1}(b),\text{proj}_{\psi_{1}(b)}(\psi_{1}(b^{\prime})))\in\partial T_{\psi_{1}(b)}.$
This application is measurable by composition. As in Lemma 4.1, the group $G$ acts on $\partial T_{1}$ via its implicit action on the chambers, using the bijection given by Proposition 2.7. With this action, $\phi$ is $G$-equivariant. Now consider
$\Phi:$ $B\times B$ $\longrightarrow T_{1}\times\partial T_{1}$
$(b,b^{\prime})$ $\longmapsto(\psi(b),\phi(b,b^{\prime}))\in T_{\psi_{1}(b)}\times\partial T_{\psi_{1}(b)},$
where $T_{1}\times\partial T_{1}$ is the metric field over $T_{1}$ defined in Lemma 4.2.
The application $\Phi$ is measurable because $\psi$ and $\phi$ are measurable. The group $G$ acts on $T_{1}\times\partial T_{1}$ via the diagonal action $g\cdot(u,\xi,C)=(gu,g\xi,gC)\in T_{gu}\times\partial T_{gu}$. Thus, for $g\in G$, and almost every $(b,b^{\prime})\in B\times B$,
$\Phi(gb,gb^{\prime})$ $=$ $(\psi(gb),\text{proj}_{\psi_{1}(gb)}(\psi_{1}(gb^{\prime})))$
[…]
$=$ $(g\psi(b),g\text{proj}_{\psi_{1}(b)}(\psi_{1}(b^{\prime}))).$
As a consequence, $\Phi$ is a $G$-map. We then have the following commutative diagram:
[…]
By Lemma 4.2, $T_{1}\times\partial T_{1}\rightarrow T_{1}$ admits a fiberwise isometric $G$-action. But $B$ is a $G$-boundary, hence the projection on the first factor $\pi_{B}:(b,b^{\prime})\in B\times B\mapsto b\in B$ is relatively isometrically ergodic. Therefore, there exists an invariant section $s:B\longrightarrow T_{1}\times\partial T_{1}$ such that $\nu\times\nu$-almost surely, $\Phi(b,b^{\prime})=s\circ\pi_{B}(b,b^{\prime})=s(b)$.
[…]
In particular, the projection $\phi(b,b^{\prime})=\text{proj}_{\psi_{1}(b)}(\psi_{1}(b^{\prime}))$ does not depend on $b^{\prime}$. We then have an essentially well defined measurable map $u:B\rightarrow X^{\infty}_{2}$ such that for almost every $b\in B$, $u(b)\in X^{\infty}_{2}$ is the unique vertex of type 2 belonging to the chamber $\text{proj}_{\psi_{1}(b)}(\psi_{1}(b^{\prime}))$. Then almost surely, $\psi_{1}(b^{\prime})$ belongs to a chamber in $\text{Res}(u(b))$, see Figure 1.
In other words, there exists a subset $\Omega_{1}$ of $B\times B$ of full $\nu\times\nu$ measure such that for all $(b,b^{\prime})\in\Omega_{1}$, $\psi_{1}(b^{\prime})$ belongs to a chamber in $\text{Res}(u(b))$. But for all $(b,b^{\prime})\in\Omega_{1}$, $\text{Res}(u(b))$ is a subset of $\partial_{\infty}X$ of circumradius $\leq\pi/2$. By Fubini, the application $\psi_{1}$ then induces an essentially well-defined $G$-map from $B$ to a subset of $\partial_{\infty}X$ of circumradius $\leq\pi/2$. By Proposition 2.3, $G$ fixes a point in $\partial_{\infty}X$, in contradiction with the assumption that the action is non-elementary.
$u(b)$$\phi(b,b^{\prime})$$\psi_{1}(b)$$\psi_{1}(b^{\prime})$
Figure 1. An apartment of the building at infinity $X^{\infty}$, in the proof of Lemma 5.3.
It remains to prove that there is no $G$-map $B\rightarrow\partial_{2}T_{1}$.
Lemma 5.4.
Under the same assumptions as in Theorem 5.1, there is no $G$-map $B\rightarrow\partial_{2}T_{1}$.
Proof.
Assume that there exists a measurable $G$-map $\psi:B\longrightarrow\partial_{2}T_{1}$. Denote by $p:\partial_{2}T_{1}\rightarrow X^{\infty}_{1}$ the natural projection, and $p\circ\psi$ by $\psi_{1}$. The application $\psi_{1}:B\rightarrow X^{\infty}_{1}$ is measurable by composition, and $G$-equivariant.
Again, we define the application $\phi$ on $B\times B$ by
$\phi:(b,b^{\prime})\in B\times B\mapsto\text{proj}_{\psi_{1}(b)}(\psi_{1}(b^{\prime}))\in\partial T_{\psi_{1}(b)}.$
As before, the application $\phi$ is $G$-equivariant, and measurable. Observe that the set $\{(b,b^{\prime})\in B\times B^{\prime}\,|\,\phi(b,b^{\prime})\in\psi(b)\}$ is $G$-invariant. By double ergodicity of $G\curvearrowright B$, its measure is full or null. Assume that almost surely, $\phi(b,b^{\prime})\notin\psi(b)$. Then we get a measurable $G$-map from $B\times B$ to the space of distinct triples of points of $\partial T_{1}$:
$\Psi:$ $B\times B\longrightarrow\partial_{3}T_{1}$
$(b,b^{\prime})\longmapsto(\psi(b),\phi(b,b^{\prime}))\in\partial_{3}T_{\psi_{1}(b)}.$
Now consider the barycenter map
[…]
$(C_{1},C_{2},C_{3})\in\partial_{3}T_{u}\longmapsto\text{bar}_{u}(C_{1},C_{2},C_{3})\in T_{u},$
By Proposition 4.3, the barycenter map just defined is measurable and $G$-equivariant, hence $\text{bar}\circ\Psi$ is a measurable $G$-map. It implies that the following diagram
[…]
is commutative. But $T_{1}\rightarrow X^{\infty}_{1}$ admits a fiberwise isometric $G$-action by Lemma 4.1. Therefore, by relative isometric ergodicity of the Poisson boundary, there exists an invariant section $s:B\longrightarrow T_{1}$, which contradicts Lemma 5.3.
As a consequence, we have that almost surely, $\phi(b,b^{\prime})\in\psi(b)$. This means that almost surely, $\phi(b,b^{\prime})$ does not depend on $b^{\prime}$. But the image of $\phi(b,b^{\prime})$ lies in $\text{Res}(\psi_{1}(b))$, which is a subset of $\partial_{\infty}X$ of circumradius $\leq\pi/2$. Now by the same arguments as for Lemma 5.3, we get a $G$-fixed point in the boundary and a contradiction. ∎
It is clear that Lemmas 5.3 and 5.4 remain true for the vertices of type 2: there is no $G$-map $B\rightarrow T_{2}$ nor $G$-maps $B\rightarrow\partial_{2}T_{2}$.
We can now prove that there exists a measurable $G$-map from $B$ to the set of chambers at infinity.
Proof of Theorem 5.1.
By assumption, $G$ acts continuously without invariant flats on the affine building $X$, and $X$ is a complete CAT(0) space of finite telescopic dimension. We can then apply Theorem 3.12: there exists a measurable $G$-map $\psi:B\rightarrow\partial_{\infty}X$.
Observe that the set $\{b\in B\;|\;\psi(b)\in\partial_{\infty}X-(X^{\infty}_{1}\sqcup X^{\infty}_{2})\}$ is measurable, and $G$-invariant. By ergodicity of $G\curvearrowright B$, its measure is null or conull. If for almost every $b\in B$, $\phi(b)$ is a regular point, then we get a $G$-map $\psi:B\rightarrow\text{ch}(X^{\infty})$, which is what we want. Otherwise, by ergodicity of the action $G\curvearrowright B$, we get either a map $\psi:B\rightarrow X^{\infty}_{1}$ or $\psi:B\rightarrow X^{\infty}_{2}$. Assume the former, the argument being identical in both cases.
We can apply Lemma 5.2: there is either a $G$-map $B\rightarrow T_{1}$, or a $G$-map $B\rightarrow\partial T_{1}$, or a $G$-map $B\rightarrow\partial_{2}T_{1}$. Along with Lemmas 5.3 and 5.4, there is a $G$-equivariant measurable map $B\rightarrow\partial T_{1}$. By Proposition 2.7, the chambers are $G$-equivariantly and measurably associated with $\partial T_{1}$, hence this gives a $G$-equivariant map $B\rightarrow\text{ch}(X^{\infty})$. ∎
5.2. Uniqueness of the boundary map
In this section, we prove that the measurable map given by Theorem 5.1 is actually unique.
Theorem 5.5.
Let $X$ be a separable complete thick building of type $\tilde{A}_{2}$ and let $G$ be a locally compact second countable group acting continuously by isometries on $X$ without invariant flats. Let $(B,\nu_{B})$ be a $G$-boundary. Then there exists a unique measurable $G$-equivariant boundary map $\psi:B\longrightarrow\text{ch}(X^{\infty})$.
Proof.
The existence of such a boundary map $\psi$ is given by Theorem 5.1. Recall that for each chamber $C\in\text{ch}({X^{\infty}})$, $C$ contains a unique vertex of type 1 and a unique vertex of type 2. Hence the boundary map $\psi:B\longrightarrow\text{ch}(X^{\infty})$ induces two $G$-maps $\psi_{1}:B\longrightarrow X^{\infty}_{1}$ and $\psi_{2}:B\longrightarrow X^{\infty}_{2}$. It is then sufficient to prove that those are the only $G$-maps $B\longrightarrow X^{\infty}_{1}$ and $B\longrightarrow X^{\infty}_{2}$. The argument being the same in both cases, we only prove that $\psi_{1}:B\longrightarrow X^{\infty}_{1}$ is unique. Assume that we have another $G$-map $\psi_{1}^{\prime}:B\longrightarrow X^{\infty}_{1}$. By ergodicity, $\psi_{1}(b)\neq\psi^{\prime}_{1}(b)$ almost surely. Since $\psi_{1}(b)$ and $\psi^{\prime}_{1}(b)$ are of type 1, there is a unique vertex of type 2 adjacent to both. Call this vertex $u(b)$. Now $\psi_{1}(b)$ and $u(b)$ belong to a unique common chamber $C_{1}(b)$, and the same goes for $\psi_{2}(b)$ and $u(b)$, whose common chamber we denote by $C_{2}(b)$. Eventually, using Proposition 2.7, this yields a $G$-map
$B$$\longrightarrow$ $\partial_{2}T_{2}$
[…]
which is in contradiction with Lemma 5.4. Hence $\psi_{1}$ is unique, proving the result. ∎
The following proposition is the analogue of a result essentially due to Furstenberg [Fur63] and Goldshei’d-Margulis [GM89], see also the exposition in Benoist-Quint [BQ16a, §9.2].
Proposition 5.6.
With the same assumptions as in Theorem 5.1, then the map $\psi_{1}=\text{proj}_{i}\circ\psi$ is essentially the unique $G$-equivariant measurable map $\psi_{i}:B\longrightarrow X^{\infty}_{i}$, for $i=1,2$.
Proof.
We are only doing the proof for $i=1$, the other case being completely similar. Let us consider $\psi^{\prime}_{1}:B\to X^{\infty}_{1}$ be a $G$-equivariant measurable map. By ergodicity of the $G$-action on $B$, the set
$\{b\in B\mid\psi_{1}(b)=\psi^{\prime}_{1}(b)\}$
is of measure 0 or 1. Let us assume by contradiction that this set is of measure 0. For (almost every) $b\in B$, let us denote by $u(b)\in X^{\infty}_{2}$ the unique vertex of type $2$ that is adjacent to both $\psi(b)$ and $\psi^{\prime}(b)$ in any apartment containing them. Alternatively, $u(b)$ can be defined as the unique vertex of type 2 adjacent to the chamber $\text{proj}_{\psi_{1}(b)}(\psi^{\prime}_{1}(b))$ as in the proof of Lemma 5.3. As we assume that almost surely $\psi_{1}(b)\neq\psi^{\prime}_{1}(b)$ we obtain that almost surely,
$\text{proj}_{u(b)}(\psi_{1}(b))\neq\text{proj}_{u(b)}(\psi^{\prime}_{1}(b)).$
Using the identification between the residue $\text{Res}(u(b))$ and the ends of the panel tree $T_{u(b)}$, we get a measurable $G$-equivariant map
$\Phi$ ::: $B\longrightarrow\partial_{2}T_{2}$
$b\mapsto(u(b),\text{proj}_{u(b)}(\psi_{1}(b)),\text{proj}_{u(b)}(\psi^{\prime}_{1}(b)))\in\partial_{2}T_{u(b)}.$
In view of Lemma 5.4, we get a contradiction. ∎
6. Stationary measure on the chambers at infinity
6.1. Uniqueness of the stationary measure
The main result of this section is the uniqueness of the stationary measure $\nu$ on the set of chambers $\text{ch}(X^{\infty})$ of the spherical building at infinity of $X$.
Theorem 6.1.
Let $X$ be a separable complete thick building of type $\tilde{A}_{2}$ and let $G$ be a locally compact second countable group acting continuously by isometries on $X$ without fixed points nor invariant flats. Let $\mu$ be an admissible symmetric measure on $G$. Then there is a unique $\mu$-stationary measure $\nu$ on $\text{ch}(X^{\infty})$. Moreover, we have the decomposition
[…]
where $\psi$ is the unique boundary map given by Theorem 5.5.
Before proving this theorem, recall the following classical result, which goes back to H. Furstenberg [Fur73, Lemma 1.33].
Lemma 6.2.
[…]
is $\mu$-stationary.
Proof of Theorem 6.1.
By Lemma 6.2, it is enough to prove that there is a unique $G$-equivariant map $B\longrightarrow\text{Prob}(\text{ch}(X^{\infty}))$. Let $\psi:B\longrightarrow\text{ch}(X^{\infty})$ be the measurable $G$-equivariant map given by Theorem 5.1. Observe that the map
$b\in B\mapsto\delta_{\psi(b)}\in\text{Prob}(\text{ch}(X^{\infty}))$
is measurable and $G$-equivariant. Let $\phi:B\rightarrow\text{Prob}(\text{ch}(X^{\infty}))$ be any $G$-equivariant map. The goal is to show that almost surely, $\phi$ is given by $\phi(b)=\delta_{\psi(b)}$.
Recall that each chamber in $X^{\infty}$ contains exactly one vertex of type 1 and one vertex of type 2, so denote by $\pi_{1}:\text{ch}(X^{\infty})\longrightarrow X^{\infty}_{1}$ and $\pi_{2}:\text{ch}(X^{\infty})\rightarrow X^{\infty}_{2}$ the corresponding maps. Let $\psi_{i}:B\rightarrow X^{\infty}_{i}$ be defined by $\psi_{i}=\pi_{i}\circ\psi:B\rightarrow X^{\infty}_{i}$, for $i=1,2$. By composition, $\psi_{i}$ is a measurable, $G$-equivariant map. Consider also the map
$\phi_{1}:b\in B\mapsto({\text{proj}_{\psi_{1}(b)}})_{\ast}\phi(b)\in\text{Prob}(\text{ch}(X^{\infty})).$
The map $\phi_{1}$ is a measurable $G$-equivariant map between $B$ and the probability measures on the chambers, but now the support of $\phi_{1}(b)$ is contained in $\text{Res}({\psi_{1}(b)})$. Now identify $\text{Res}(\psi_{1}(b))$ with the ends $\partial T_{\psi_{1}(b)}$ of the panel tree at $\psi_{1}(b)$ given by Proposition 2.7. Therefore, $\phi_{1}$ can be written as
$\phi_{1}:B$$\rightarrow$ $\text{Prob}(\partial T_{1})$
$b$ $\mapsto$ $\phi_{1}(b)\in\text{Prob}(\partial T_{\psi_{1}(b)}).$
By ergodicity of the Poisson boundary, the cardinal $k(b)$ of the support of $\phi_{1}(b)$ is almost surely constant. If almost surely $k(b)\geq 3$, then the barycenter map constructed in Proposition 4.6 associates to the support of $\phi(b)$ a canonical point in $T_{\psi_{1}(b)}$. This gives a measurable $G$-equivariant map $B\longrightarrow T_{1}$, which is impossible due to Lemma 5.3.
If the support of the measure $\phi_{1}(b)$ is almost surely of cardinal $2$, then we have a $G$-map $B\longrightarrow\partial_{2}T_{1}$, which is in contradiction with Lemma 5.4.
Then $\phi_{1}(b)$ has to be a Dirac mass, which we denote by $\delta_{\xi(\omega)}$ for $(\psi_{1}(\omega),\xi(\omega))\in\partial T_{\psi_{1}(\omega)}$. Again, by Proposition 2.7, this gives a $G$-map $B\longrightarrow\text{ch}(X^{\infty})$. But by Theorem 5.5, any such map is unique and equal to $\psi$. Therefore, $\phi_{1}(b)$ is the Dirac mass at $\psi(b)$.
What we obtained so far is that for any $C\in\text{ch}(X^{\infty})$,
(7) $\phi_{1}(b)(\text{proj}_{\psi_{1}(b)}(C))$ $=$ $\delta_{\psi(b)}(\text{proj}_{\psi_{1}(b)}(C)).$
By definition, $\phi_{1}(b)=({\text{proj}_{\psi_{1}(b)}})_{\ast}\phi(b)$. Then by Equation (7), the support of $\phi(b)$ is contained in the set of chambers $C\in\text{ch}(X^{\infty})$ such that $\text{proj}_{\psi_{1}(b)}(C)=\psi(b)$.
By repeating the same argument as before for $\psi_{2}:B\rightarrow X^{\infty}_{2}$, we obtain that $\phi(b)$ must also be supported on the set of chambers $C\in\text{ch}(X^{\infty})$ such that $\text{proj}_{\psi_{2}(b)}(C)=\psi(b)$. Combining these two results, we get that $\phi(b)$-almost surely, $\text{proj}_{\psi_{1}(b)}(C)=\text{proj}_{\psi_{2}(b)}(C)=\psi(b)$, hence $\phi(b)$ is supported on the chamber $\psi(b)$, meaning that $\phi(b)$ is essentially equal to $\delta_{\psi(b)}$. This proves the uniqueness of $\psi$, therefore of the stationary measure. ∎
Let us denote by $\{Z_{n}\}$ the random walk generated by $\mu$ on $G$, as in the assumptions of Theorem 6.1, and denote by $B$ the Poisson-boundary of $(G,\mu)$. An immediate application of Theorem 3.3 gives the following useful result.
Corollary 6.3.
Let $\nu$ be the unique $\mu$-stationary measure on $\text{ch}(X^{\infty})$ associated to $\mu$, and let $\psi:B\to\text{ch}(X^{\infty})$ be the unique boundary map between $B$ and the chambers at infinity given by Theorem 5.5. Then $\mathbb{P}$-almost surely, $Z_{n}(\omega)\nu\rightarrow\delta_{\psi(\text{bnd}(\omega))}$ in the weak-$\ast$ topology.
6.2. Opposite chambers at infinity
The following is an important feature of the stationary measure $\nu$. Recall that two chambers $C,C^{\prime}$ in $X^{\infty}$ are opposite if they are at maximal gallery distance, i.e. if $d_{\Gamma}(C,C^{\prime})=3$. Since apartments are convex, the convex hull of two chambers is contained in any apartment containing both of them. As a consequence, two opposite chambers are contained in a unique apartment. Denote by $\nu$ the unique $\mu$-stationary measure on $\text{ch}(X^{\infty})$ given by Theorem 6.1.
Proposition 6.4.
Let $\nu$ be the unique $\mu$-stationary measure on $\text{ch}(X^{\infty})$ associated to $\mu$, as given by Theorem 6.1. Then $\nu\otimes\nu$-almost every pair of chambers in $\text{ch}(X^{\infty})$ are opposite.
Proof.
Consider the $G$-equivariant map $\psi:B\rightarrow\text{ch}(X^{\infty})$ given by Theorem 5.1, and the following measurable map :
ϕ::italic-$\phi:$ $B\times B$ $\longrightarrow W_{0}$
$(b,b^{\prime})$ $\longmapsto\delta(\psi(b),\psi(b^{\prime})),$
where $\delta$ is the Weyl distance function associated to the spherical Coxeter group $(W_{0},S)$ of type $A_{2}$, associated to the spherical building at infinity $X^{\infty}$. Then by $G$-equivariance of $\psi$, $\phi$ is a $G$-invariant measurable map. By double ergodicity of the Poisson boundary, $\phi$ is then essentially constant.
If its essential value is $e$, then almost surely, $\psi(b)=\psi(b^{\prime})$, and there is a $G$-fixed chamber on the boundary, which is impossible because the action is non-elementary.
Let us denote by $l_{S}$ the word metric on $W$ given by the set $S$. From now on, let $B_{0},B_{1}\subseteq B$ be conull sets such that for all $(b,b^{\prime})\in B_{0}\times B_{1}$, $\phi(b,b^{\prime})$ is the essential value of $\phi$. If almost surely, $l_{S}(\phi(b,b^{\prime}))=1$, then almost surely $\psi(b)$ and $\psi(b^{\prime})$ are adjacent. Since $\phi$ is constant, then almost surely $\psi(b)$ and $\psi(b^{\prime})$ contain a vertex of constant type. Let $b\in B_{0}$, denote by $u(b)\in\psi(b)$ the unique vertex of type $1$. Then for all $b^{\prime}\in B_{1}$, $\psi(b^{\prime})\in\text{Res}(u(b))$. Again, because the action is non-elementary, this is impossible by Proposition 2.3 since $\text{Res}(u(b))$ has radius $\leq\pi/2$.
If $l_{S}(\phi)$ is almost surely equal to 2. Then the gallery between $\psi(b)$ and $\psi(b^{\prime})$ is of constant type $(1,2)$ or $(2,1)$. But this is impossible because $\delta(\psi(b^{\prime}),\psi(b))=\delta(\psi(b),\psi(b^{\prime}))^{-1}$, so by ergodicity, we have that $\delta(\psi(b),\psi(b^{\prime}))=\delta(\psi(b),\psi(b^{\prime}))^{-1}$.
Consequently, $\psi(b)$ and $\psi(b^{\prime})$ are almost surely opposite. The proposition then follows from the decomposition
[…]
given by Theorem 6.1. ∎
7. Convergence to a chamber at infinity
Let $X$ be a separable complete building of type $\tilde{A}_{2}$ and let $G$ be a locally compact second countable group acting continuously by isometries and non-elementarily on $X$. Let $\mu$ be an admissible measure on $G$. As in Section 3, we let $(\Omega,\mathbb{P})$ be the probability space $(G^{\mathbb{N}},\mu^{\otimes\mathbb{N}})$, with the product $\sigma$-algebra, and $\{Z_{n}(\omega)\}$ be the random walk on $G$ associated to $\mu$. The goal of this section is to prove that for any basepoint $o\in X$, the random walk $(Z_{n}o)$ converges almost surely to a chamber at infinity. In fact, under the assumption that $\mu$ has finite second moment, we prove that the Lyapunov spectrum of the random walk is simple.
7.1. Regular sequences in buildings
In this section, we present the notion of regular sequences and relate it with the notion of Lyapunov regularity. This notion first appeared in [Ki87] for symmetric spaces, and then developed in [PW15] for buildings.
We fix $\mathfrak{a}^{+}$ a fundamental closed Weyl chamber for the affine reflection group $(W,V)$. For any two points $x,y\in X$, the type $\theta(x,y)$ is a vector in $\mathfrak{a}^{+}$.
Definition 7.1.
Let $\lambda\in\mathfrak{a}^{+}$ be a fixed vector. A embedding $\gamma:\mathbb{R}_{+}\to X$ is called a $\lambda$-regular ray if for any $s,t\in\mathbb{R}_{+}$,
$\theta(\gamma(t),\gamma(t+s))=s\lambda\in\mathfrak{a}^{+}.$
A sequence $(x_{n})$ in $X$ is called $\lambda$-regular if there exists a $\lambda$-regular ray $\gamma$ such that
$d(x_{n},\gamma(n))=o(n).$
The following theorem is the main result of [PW15], where the authors only consider discrete buildings. In the context of symmetric spaces, it was proven by Kaimanovich in [Ki87, Theorems 2.1 and 2.4]
Theorem 7.2.
Let $(x_{n})$ be a sequence in an affine building $X$, and let $\lambda\in\mathfrak{a}^{+}$. Then the following statements are equivalent:
  1. (1)
    the sequence $(x_{n})$ is $\lambda$-regular;
  2. (2)
    $d(x_{n},x_{n+1})=o(n)$ and
    $\frac{1}{n}\theta(o,x_{n})\longrightarrow\lambda.$
Most of the proof in [PW15, Theorem 3.2] goes through to non-discrete buildings. We denote by $\{\alpha_{i}\}_{i}$ the set of roots, and by $\{\lambda_{i}\}_{i\in I_{0}}$ be the basis of $V$ dual of $\{\alpha_{i}\}_{i}$, so that $\langle\alpha_{i},\lambda_{j}\rangle=\delta_{i,j}$, which we call coweights. The “only if” part is straightforward.
Proof.
(1 $\Rightarrow$ 2) Let us first assume that $(x_{n})$ is $\lambda$-regular. Let $\gamma:[0,\infty[\rightarrow X$ be a $\lambda$-ray. If $\lambda=0$, the result is immediate. We can then assume without loss of generality that $\lambda\neq 0$ and $o=\gamma(0)$. Then by the triangular inequality,
$d(x_{n},x_{n+1})$$\leq$ $d(x_{n},\gamma(n))+d(\gamma(n),\gamma(n+1))+d(\gamma(n+1),x_{n+1})=o(n).$
Since $\gamma$ is a $\lambda$-ray starting at $o$, $\theta(o,\gamma(n))=n\lambda$ for all $n$. Now
$\theta(o,x_{n})=\sum_{i}\langle\theta(o,x_{n}),\alpha_{i}\rangle\lambda_{i}.$
By he law of cosines in $\text{CAT}(0)$ spaces (see for instance [BH99, Exercise II.1.9]),
$d(x_{n},\gamma(n))^{2}\geq d(o,x_{n})^{2}+d(o,\gamma(n))^{2}-2d(o,x_{n})d(o,\gamma(n))\cos(\angle_{o}(x_{n},\gamma(n))).$
Since $d(x_{n},\gamma(n))=o(n)$ and $d(o,\gamma(n))=n\|\lambda\|$, we obtain that $\angle_{o}(x_{n},\gamma(n))\rightarrow_{n}0$. AS a consequence, for all $i\in I_{0}$ we have that
$\langle\theta(o,x_{n}),\alpha_{i}\rangle=\langle\theta(o,\gamma(n)),\alpha_{i}\rangle+o(n),$
whence the result. ∎
(2 $\Rightarrow$ 1) The approach of this part given in [PW15] goes through without modifications, as it mostly relies on $\text{CAT}(0)$ geometry and Lemma 7.3 below, whose proof in non-discrete buildings is exactly the same, see also [Rou23, 2.4.11].
Lemma 7.3.
Let $o,x,y\in X$ be such that $\theta(o,x)=\theta(o,y)=\lambda$ and for which $[o,x]\cap[o,y]=\{o\}$. Then there is $C>0$ depending only on the direction of $\lambda$ (not on its length) such that $d(x,y)\geq Cd(o,x)$. Actually,
$C=\min\{2\sin(\angle(\lambda,w\lambda))\mid w\in W_{0}-\{1\}\}.$
7.2. Convergence of the random walk
We keep the notations introduced at the beginning of the section. In particular, we assume the measure $\mu$ on $G$ to be symmetric, and to have finite second moment
[…]
The goal of this section is to show that for $o\in X$, the $\mu$-random walk $(Z_{n}o)$ converges to a regular point of the visual boundary.
Recall that the escape rate (or drift) of the random walk is defined as
[…]
if $\mu$ has finite first moment, and $l_{X}(\mu)=\infty$ otherwise.
We are going to use the following theorem from H. Izeki.
Theorem 7.4 ([Ize23, Theorem A]).
Let $Y$ be a complete CAT(0) space which is either proper or of finite telescopic dimension, and let $G$ be a discrete countable group equipped with a symmetric and admissible probability measure $\mu$ with finite second moment. Assume that $G\curvearrowright Y$ is a non-elementary action. Then the drift of the random walk is positive:
$l_{Y}(\mu)>0.$
In our case, we consider a metrically complete $\tilde{A}_{2}$-building, hence of finite telescopic dimension (equal to 2), and the theorem applies. We will combine this with the celebrated result by Karlsson and Margulis [KM99, Theorem 2.1] on convergence of ergodic cocycles in non-positive curvature.
The main result of this section is the following theorem.
Theorem 7.5 (Simplicity of the Lyapunov spectrum).
Let $X$ be a separable complete building of type $\tilde{A}_{2}$ and let $G$ be a locally compact second countable group acting continuously by isometries and non-elementarily on $X$. Let $\mu$ be an admissible measure on $G$ with finite second moment. Then there exists a regular $\lambda\in\mathfrak{a}^{++}$ such that for almost every $\omega\in\Omega$, $(Z_{n}(\omega))$ is $\lambda$-regular. In particular, $(Z_{n}o)$ converges almost surely to a regular point of the visual boundary.
Recall that if we denote by $(B,\nu_{B})$ the Poisson-Furstenberg boundary associated to $(G,\mu)$, we have a canonical measurable map
$\text{bnd}:\Omega\rightarrow B$
such that $\text{bnd}_{\ast}\mathbb{P}=\nu_{B}$.
Proof.
By Theorem 7.4, the drift $l_{\mu}:=l_{X}(\mu)$ of the random walk $(Z_{n}(\omega)o)$ is positive. We can then apply [KM99, Theorem 2.1]: for almost every $\omega$, there exists a geodesic ray $\gamma^{\omega}$ such that
$\lim\frac{1}{n}d(\gamma^{\omega}(l_{\mu}n),Z_{n}(\omega)o)=0.$
In other words, if we denote $\lambda^{\omega}=\theta(\gamma^{\omega}(0),\gamma^{\omega}(l_{\mu}))\in\mathfrak{a}^{+}$ whenever it is defined, then $(Z_{n}(\omega)o)$ is almost surely $\lambda^{\omega}$-regular. In particular, the random walk converges almost surely to a point of the visual boundary. We then obtain a $G$-equivariant measurable map
$F:B\to\partial_{\infty}X,$
which to almost every $b\in B$ associates the limit point of $(Z_{n}(\omega))$ for $\omega\in\text{bnd}^{-1}(b)$. By ergodicity of the $G$-action on the Poisson boundary, this boundary point is either almost surely a regular point, a vertex at infinity of type 1 or a vertex at infinity of type 2. Note that this decomposition corresponds exactly to the type of $\lambda_{\mu}$: $\lambda_{\mu}$ is regular (resp. of type 1,2) if and only if $(Z_{n}(\omega))$ converges to a regular point of the boundary (resp. to a vertex of type 1,2).
Now consider $\check{\mu}=i_{\ast}\mu$ for $i(g)=g^{-1}$. The measure $\mu$ being symmetric by assumption, we have $\lambda_{\mu}=\lambda_{\check{\mu}}$. However, if we denote by $\iota:\mathfrak{a}^{+}\to\mathfrak{a}^{+}$ the involution, we have that $\lambda_{\check{\mu}}=\iota(\lambda_{\mu})$, see for instance [BQ16a, Corollary 9.11]. In $\tilde{A}_{2}$ buildings, the involution changes the type. In particular, the Lyapunov vector $\lambda_{\mu}$, being involution-invariant, is regular. ∎
The following corollary states that in particular, whenever $\mu$ has finite second moment, the random walk converges to a chamber at infinity in the polyhedral sense too. For a detailed account on this notion, we refer to [CL11] (for discrete buildings) and to [Rou23, §3.4].
Corollary 7.6.
We keep the assumption of Theorem 7.5. Let $o,x\in X$. Then for almost every $\omega$, there exists $n_{0}$ such that for all $n\geq n_{0}$, the projection $\Sigma_{o}(Z_{n}(\omega)x)$ of $Z_{n}(\omega)x$ on the residue building $\Sigma_{o}X$ is the chamber $\Sigma_{o}(\tilde{\psi}(\omega))$, where $\tilde{\psi}(\omega)=\psi(\text{bnd}(\omega))\in\text{ch}(X^{\infty})$ is the boundary map given by Theorem 5.1.
Proof.
By Theorem 7.5, there exists $\lambda\in\mathfrak{a}^{++}$ a regular vector and $\Omega^{\prime}\subseteq\Omega$ of full measure satisfying: for all $\omega\in\Omega^{\prime}$, there exists a $\lambda$-ray $\gamma^{\omega}$ such that the random walk $(Z_{n}(\omega)x)$ sublinearly tracks $\gamma^{\omega}$. Fix $\omega\in\Omega^{\prime}$, and write $\gamma:=\gamma^{\omega}$. We can assume without loss of generality that $\gamma(0)=o$, so that $\gamma$ is contained in the interior of the Weyl chamber $S=Q(o,C_{\omega})$, where $C_{\omega}=\tilde{\psi}(\omega)$. The vector $\lambda$ is regular, hence the distance between $\gamma(n)$ and any of the walls delimiting $S$ grows linearly. Since $d(x_{n},\gamma(n))=o(n)$, there exists $N$ such that for all $n\geq N$, $d(x_{n},\gamma(n))$ is strictly less than the distance from $\gamma(n)$ and any of these walls. Let $n\geq N$. We need to prove that the germ of the geodesic segment $[o,x_{n}]$ is contained in $\text{germ}(S)$. Let $S_{n}$ be a Weyl chamber based at $o$ containing the $[o,x_{n}]$. By [Par00, Proposition 1.15], there exists an apartment $A_{n}$ containing $S$ and the germ of $S_{n}$ in $o$. Now consider the retraction $\rho_{A_{n},S}$ onto $A_{n}$ centered on $S$. The retraction does not increase the distance
$d(\rho_{A_{n},S}(x_{n}),\rho_{A_{n},S}(\gamma(n)))=d(\rho_{A_{n},S}(x_{n}),\gamma(n))\leq d(x_{n},\gamma(n)),$
and as $n\geq N$, $\rho_{A_{n},S}(x_{n})$ is contained in the interior of $S$. As $\rho_{A_{n},S}$ induces a retraction of $\Sigma_{o}X$ onto $\Sigma_{o}A_{n}$, we obtain that the germ of $[o,x_{n}]$ is strictly contained in the $\text{germ}_{o}(S)$ and the result follows. ∎
7.3. Stationary measures on $X\cup\text{ch}(X^{\infty})$
In this subsection, we show that the measure on $\text{ch}(X^{\infty})$ given by Theorem 6.1 is actually the only stationary measure on the bordification $X\cup\text{ch}(X^{\infty})$. This is a straightforward consequence of the convergence of the random walk to the visual boundary described by Theorem 7.5.
Let us then endow $X\cup\text{ch}(X^{\infty})$ with the topology on $X\cup\text{ch}(X^{\infty})$ described in Section 2. This gives a metrizable separable Hausdorff topology on $X\cup\text{ch}(X^{\infty})$.
Proposition 7.7.
We keep the notations of Theorem 7.5. Let $\nu$ be any $\mu$-stationary probability measure supported on the bordification $X\cup\text{ch}(X^{\infty})$ for the cone topology. Then the support of $\nu$ is in $\text{ch}(X^{\infty})$.
Proof.
Let $o\in X$. Up to decomposing the measure $\nu$, we assume by contradiction that $\nu$ is supported in $X$. Then there exists $r>0$ such that the open ball $O:=B(o,r)$ satisfies $\nu(O)\geq\frac{2}{3}$. Since $\nu$ is $\mu$-stationary, we can apply Theorem 3.3: there exists an essentially well-defined $G$-equivariant measurable map $\omega\mapsto\nu_{\omega}$ such that $Z_{n}(\omega)\nu\rightarrow\nu_{\omega}$ almost surely in the weak-$\ast$ topology. and such that we have the decomposition
[…]
In view of Theorem 6.1, we derive the following.
Corollary 7.8.
We keep the notations of Theorem 7.5. There is a unique $\mu$-stationary probability measure $\nu$ on $X\cup\text{ch}(X^{\infty})$, and $\nu$ is given by Theorem 6.1.